### v1

# Asymptotic constraints for 1D planar grey photon diffusion from linear transport with special-relativistic effects

Abstract
We derive a grey linear diffusion equation for photons with respect to inertial (or lab-frame) space and time, using asymptotic analysis in 1D planar geometry. The solution of the equation is the comoving radiation energy density. Our analysis does not make use of assumptions about the magnitude of velocity; instead we derive an asymptotic scaling in the lab frame such that we avoid apparent non-physical pathologies that are encountered with the standard static-matter scaling. We permit the photon direction to be continuous (as opposed to constraining the analysis to discrete ordinates). The result is a drift-diffusion equation in the lab frame for comoving radiation energy density, with an adiabatic term that matches the standard semi-relativistic diffusion equation. Following a recent study for discrete directions, this equation reduces to a pure advection equation as the velocity approaches the speed of light. We perform preliminary numerical experiments comparing solutions to relativistic lab-frame Monte Carlo transport and to the well-known semi-relativistic diffusion equation.
1 Introduction
Radiation (photon, neutrino) diffusion through optically thick matter moving at (near-)relativistic speed or in the presence of gravity effects is a ubiquitous phenomenon in the study of astrophysical plasmas. It is well known to be a limiting behavior of the more general phenomenon of radiative transfer. For instance, photon propagation through expanding supernova ejecta can be characterized with a diffusion solution (see, for instance, [arnett1982, pinto2000]). Numerical implementation of the diffusion equation can be used to accelerate radiative transfer in optically thick regions of space and frequency. One example is Monte Carlo (MC) accelerated with random walk (RW) or discrete diffusion Monte Carlo (DDMC) [fleck1984, gentile2001, densmore2007, densmore2012, cleveland2014, keady2017, smith2020]. These RW and DDMC methods have been developed for applications to: neutrinos in core collapse supernova engines [abdikamalov2012], cosmic rays [harding2016, fitzaxen2021], radiation pressure feedback in stellar cluster dynamics [tsang2018], supernova and kilonova spectral synthesis [wollaeger2013, wollaeger2014, wagle2023], and Lyman $\alpha$ photon transfer during galaxy evolution [smith2018].
While diffusion theory has found practical use in describing photon and neutrino fields in astrophysical plasmas, the infinite signal speed of the diffusion equation is incompatible with special relativity (see, for instance, [dunkel2007, gzyl2025]). This can readily be seen for static material backgrounds, where the Green’s function conforms to a Gaussian distribution, which has non-zero values everywhere. As a result, the applicability of diffusion is limited to high-optical depth regions, unless the equation is modified to restrict the signal speed (for instance $P_{1}$, $M_{1}$ [levermore1984] and flux-limit diffusion [levermore1981]). We may also see the effect of Lorentz transforming the space-time derivatives in the diffusion equation. Suppose that a frame of reference, denoted by subscript 0, possesses particles that adhere to a 1D diffusion equation,
$\frac{\partial\phi_{0}}{\partial t_{0}}-D\frac{\partial^{2}\phi_{0}}{\partial x_{0}^{2}}=0\;\;,$
where $t_{0}$ is time, $x_{0}$ is the spatial coordinate, $D$ is a constant diffusion coefficient, and $\phi_{0}$ is the particle density. Applying the standard Lorentz transformations for the partial derivatives (see, for instance, [lowrie2023]) gives
$\gamma\left(\frac{\partial}{\partial t}+v\frac{\partial}{\partial x}\right)\phi_{0}-D\left[\gamma\left(\frac{\partial}{\partial x}+\frac{v}{c^{2}}\frac{\partial}{\partial t}\right)\right]^{2}\phi_{0}=0\;\;,$
where $v$ is some velocity along $x$, $c$ is the speed of light in vacuum, $\gamma=1/\sqrt{1-\beta^{2}}$ and $\beta=v/c$. Assuming $v$ is constant in $x$ and $t$, this equation can be expanded to
$\left(\frac{\partial}{\partial t}+c\beta\frac{\partial}{\partial x}\right)\phi_{0}-\gamma D\left(\frac{\partial^{2}}{\partial x^{2}}+2\frac{\beta}{c}\frac{\partial^{2}}{\partial t\partial x}+\frac{\beta^{2}}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}\right)\phi_{0}=0\;\;.$
Thus in the new frame, distinguished here by the absence of a subscript, a second derivative in time is manifest. Similar to the infinitely extended signal in space, this second derivative in time must couple all time from the past to the future and vice versa. Consequently, this equation is not immediately amenable to numerical schemes that can only advance forward in time in the frame of $(x,t)$.
The problem of unifying diffusion with the principles of special relativity has been recognized in some form for approximately one century; some approaches to combining them include: (i) statistical analysis of stochastic kinematics and (ii) asymptotic analysis of the radiative transfer equation with relativistic corrections. The latter approach is often pursued in transport theory and astrophysics literature, where the effects of collective fluid motion on particle distributions is a focus.
Stochastic analyses originated by [hakim1965] and [dudley1966] (and further developed in subsequent works) demonstrate that relativity breaks Markovian processes in space, where one would expect that the set of spatial positions for the current particle state completely determine the next state [dudley1966, dunkel2007, giona2017]. In developing the relativistic theory, one might first consider bounding the particle speed in a Poisson-Kac process [kac1974], which corresponds to the Telegrapher ($P_{1}$) equation and is non-Markovian [kac1974, dunkel2007]. Spatially non-Markovian stochastic processes have been formulated in special [dunkel2007, serva2020] and general [herrmann2010] relativistic settings. Properly accounting for relativity in a collective sense, where a frame is moving relative to the frame in which diffusion is a good description, is a somewhat distinct problem. To this end, [giona2017] examine the particle motion as a Poisson-Kac process along discrete directions, where the process exists in its usual form in one inertial (comoving) frame, finding in 1D that the effective diffusion coefficient in another inertial frame is $1/\gamma^{3}$ times the comoving coefficient; [giona2017] observe that this follows the intuition that $D\sim\delta x^{2}/\delta t\sim(\delta x_{0}/\gamma)^{2}/(\gamma\delta t_{0})$, from Lorentz contraction of space and time dilation.
For asymptotic analysis, the principal assumption is typically that the inverse of the macroscopic opacity is large compared to all length scales of interest [thomas1930]. Unlike the works above that focus on the statistical mechanics of the particles, this approach usually: starts from the full radiative transfer equation, imposes the assumed parameter scalings, and expands the radiation intensity in powers of a small-scale parameter (sometimes labeled $\varepsilon$). Asymptotic analysis has been applied to deriving a semi-relativistic diffusion equation [morel2006, lowrie2014], where $\beta$ is assumed to be asymptotically small (equivalently: the hydrodynamic flow time scale is assumed much longer than the light crossing time at the length scale of interest). In the fully relativistic context, including general relativity, asymptotic analysis has been examined early on by [thomas1930] and subsequently generalized by [anderson1972] (see also [achterberg2018]). Building off of these works, [shibata2011] demonstrate the truncated Thorne moment formalism (see [thorne1981]) converges to a diffusion equation with modifications from fluid acceleration and shear in the optically thick limit. The analyses of [anderson1972] and [shibata2011] consider both asymptotically large absorption and scattering opacity. This is consistent with the static-material scaling that furnishes equilibrium diffusion from non-linear thermal radiative transfer, but departs in not assuming that the light crossing distance on the time scale of interest is large compared to the spatial domain of interest [larsen1983].
In the present study, we instead consider an asymptotic analysis that restricts to a special relativistic, 1D-planar, linear transport context, and we focus on the correspondence of the lab frame (some inertial frame) to the non-inertial frame comoving with fluid parcels of the background matter. In a departure from the work of [anderson1972] and [shibata2011] we attempt to use the standard parameter scaling for linear transport (see, for instance, [habetler1975, larsen1992]) in the comoving frame. Thus the scattering opacity is assumed to be asymptotically large while the absorption opacity is assumed to be asymptotically small, and there is not an equilibrium solution (for instance, the Planck or Fermi-Dirac distributions in the comoving frame). In doing so, our objective is to: (i) highlight apparent non-physical pathologies that occur from applying this standard scaling when velocity gradients are present, (ii) demonstrate that a particular scaling furnishes the Poisson-Kac-derived $\gamma$-dressed result of [giona2017], but for continuous directions and including the standard grey contribution from adiabatic expansion and Doppler shift (see, for instance [castor2004]), (iii) compare solutions of the diffusion equation to a relativistic linear transport solution and better-known semi-relativistic diffusion equation [castor2004].
This paper is organized as follows. In Section 2, we present all considerations for deriving the fully relativistic 1D planar grey diffusion equation, starting from a linear transport equation with isotropic opacity and elastic isotropic scattering in the comoving frame. This includes considering the effect of frequency Doppler shift on integration over frequency for the lab-frame grey equation, in Section 2.1. We present properties of a simple class of functions, which are weighted powers of Lorentz transform (Doppler shift) factors of frequency, and their associated integrals in Section 2.2. These functions and their properties facilitate both the asymptotic derivation and the lab-frame harmonic expansion (Appendix AB). In Section 2.4, we present the development of the lab-frame asymptotic scaling that furnishes a closed comoving equation with respect to lab-frame space-time coordinates. This section is subdivided into sections that showcase the pathologies encountered from simpler asymptotic scaling attempts. In Section 3, we present a discretization of the fully relativistic diffusion equation, identifying terms with corresponding non- or semi-relativistic DDMC analogs. We then compare this discretization to lab-frame MC transport to provide a preliminary exploration of where the fully relativistic diffusion equation is applicable in Section 4. In the appendix, we provide: (A) details of the integrated Doppler function recursion, (B) the aforementioned harmonic ($P_{1}$) expansion, and finally (C) a brief non-relativistic attempt of the apparent opacity-only scaling of [thomas1930, anderson1972, shibata2011].
2 Derivation of 1D planar fully relativistic grey diffusion
2.1 Grey transport with isotropic comoving scattering
Assuming 1D planar, grey, linear transport in the inertial lab frame,
$\frac{1}{c}\frac{\partial\psi}{\partial t}+\mu\frac{\partial\psi}{\partial x}+\sigma_{t}\psi=\int_{-1}^{1}R_{s}(\mu^{\prime}\rightarrow\mu)\psi^{\prime}d\mu^{\prime}+\frac{q}{2}\;\;,$ (1)
where $c$ is the speed of light, $t$ is time, $x$ is position, $\mu$ is the x-component of direction, $\sigma_{t}$ is the lab-frame total opacity, $R_{s}(\mu^{\prime}\rightarrow\mu)$ is the scattering redistribution kernel for angle, $q$ is a lab-frame source term, and $\psi$ is the lab-frame intensity. In subsequent sections, we will also use
$\phi=cE=\int_{-1}^{1}\psi\,d\mu\;\;,$ (2a)
$F=\int_{-1}^{1}\mu\psi\,d\mu\;\;,$ (2b)
$P=\int_{-1}^{1}\mu^{2}\psi\,d\mu\;\;,$ (2c)
where $\psi$, $E$, $F$ and $P$ are the scalar intensity, energy density, flux, and pressure of the radiation field. Comoving frame versions of all defined quantities will be subscripted with 0.
We first derive the lab-frame scattering kernel under the assumption of frequency-independent, isotropic, elastic scattering in the comoving frame. Frequency and the full direction vector, $\hat{\Omega}=(\mu,\eta,\xi)$, are included to more readily use Lorentz transform invariants. Neglecting spatial gradients and sources, the comoving frame equation is
$\sigma_{t,0}\psi_{0}(\nu_{0},\hat{\Omega}_{0})=\sigma_{s,0}\int_{4\pi}\int_{0}^{\infty}\frac{\delta(\nu_{0}^{\prime}-\nu_{0})}{4\pi}\psi_{0}(\nu_{0}^{\prime},\hat{\Omega}_{0}^{\prime})d\nu_{0}^{\prime}d\Omega_{0}^{\prime}\;\;,$ (3)
where $\delta(\cdot)$ is the Dirac delta distribution. If the opacity is purely scattering, $\sigma_{t,0}=\sigma_{s,0}$, integrating Eq. (3) over solid angle would give $0=0$, indicating conservation. Making use of the Lorentz invariants and transformation,
$\left(\frac{\nu_{0}}{\nu}\right)^{2}\sigma_{t}\psi(\nu,\hat{\Omega})=\frac{1}{4\pi}\sigma_{s,0}\int_{4\pi}\int_{0}^{\infty}\delta\left(\nu^{\prime}\gamma(1-\hat{\Omega}^{\prime}\cdot\vec{\beta})-\nu\gamma(1-\hat{\Omega}\cdot\vec{\beta})\right)\left(\frac{\nu_{0}^{\prime}}{\nu^{\prime}}\right)^{3}\psi(\nu^{\prime},\hat{\Omega}^{\prime})\left(\frac{\nu^{\prime}}{\nu_{0}^{\prime}}\right)d\nu^{\prime}d\Omega^{\prime}\;\;,$ (4)
[…]
$\left(\gamma(1-\vec{\beta}\cdot\hat{\Omega})\right)^{2}\sigma_{t}\psi(\nu,\hat{\Omega})=\\ \frac{1}{4\pi}\sigma_{s,0}\int_{4\pi}\left(\gamma(1-\vec{\beta}\cdot\hat{\Omega}^{\prime})\right)^{2}\left\{\int_{0}^{\infty}\delta\left(\nu^{\prime}\gamma(1-\hat{\Omega}^{\prime}\cdot\vec{\beta})-\nu\gamma(1-\hat{\Omega}\cdot\vec{\beta})\right)\psi(\nu^{\prime},\hat{\Omega}^{\prime})d\nu^{\prime}\right\}d\Omega^{\prime}\;\;.$ (5)
Using the following substitution pattern,
$\int_{0}^{\infty}\delta(A\nu^{\prime}-B\nu)f(\nu^{\prime})d\nu^{\prime}=\frac{1}{A}\int_{0}^{\infty}\delta(u^{\prime}-B\nu)f(u^{\prime}/A)du^{\prime}=\frac{1}{A}f\left(\frac{B}{A}\nu\right)$
the inner integral over pre-scatter lab-frame frequency can be simplified,
$\left(\gamma(1-\vec{\beta}\cdot\hat{\Omega})\right)^{2}\sigma_{t}\psi(\nu,\hat{\Omega})=\frac{1}{4\pi}\sigma_{s,0}\int_{4\pi}\left(\gamma(1-\vec{\beta}\cdot\hat{\Omega}^{\prime})\right)^{2}\left\{\frac{1}{\gamma(1-\vec{\beta}\cdot\hat{\Omega}^{\prime})}\psi\left(\left(\frac{1-\vec{\beta}\cdot\hat{\Omega}}{1-\vec{\beta}\cdot\hat{\Omega}^{\prime}}\right)\nu,\hat{\Omega}^{\prime}\right)\right\}d\Omega^{\prime}\\ =\frac{\gamma}{4\pi}\sigma_{s,0}\int_{4\pi}(1-\vec{\beta}\cdot\hat{\Omega}^{\prime})\psi\left(\left(\frac{1-\vec{\beta}\cdot\hat{\Omega}}{1-\vec{\beta}\cdot\hat{\Omega}^{\prime}}\right)\nu,\hat{\Omega}^{\prime}\right)d\Omega^{\prime}\;\;.$ (6)
Integrating Eq. (6) over lab-frame frequency,
$\left(\gamma(1-\vec{\beta}\cdot\hat{\Omega})\right)^{2}\sigma_{t}\psi(\hat{\Omega})=\frac{\gamma}{4\pi}\sigma_{s,0}\int_{4\pi}(1-\vec{\beta}\cdot\hat{\Omega}^{\prime})\left(\frac{1-\vec{\beta}\cdot\hat{\Omega}^{\prime}}{1-\vec{\beta}\cdot\hat{\Omega}}\right)\psi(\hat{\Omega}^{\prime})d\Omega^{\prime}\;\;,$ (7)
where we have dropped the $\nu$ argument in $\psi$,
$\psi(\hat{\Omega})=\int_{0}^{\infty}\psi(\nu,\hat{\Omega})d\nu\;\;.$
The factor multiplying $\psi(\hat{\Omega}^{\prime})$ comes from the following substitution,
$\int_{0}^{\infty}f\left(\frac{B}{A}\nu\right)d\nu=\frac{A}{B}\int_{0}^{\infty}f(u^{\prime})du^{\prime}\;\;.$
Further simplifying Eq. (7),
$\sigma_{t}\psi(\hat{\Omega})=\frac{1}{4\pi}\left(\frac{\sigma_{s,0}}{\gamma(1-\vec{\beta}\cdot\hat{\Omega})^{3}}\right)\int_{4\pi}(1-\vec{\beta}\cdot\hat{\Omega}^{\prime})^{2}\psi(\hat{\Omega}^{\prime})d\Omega^{\prime}\;\;.$ (8)
Assuming velocity $\vec{\beta}=\beta\hat{x}$, where $\hat{x}$ is the unit vector in the $x$-direction, and $\psi$ is symmetric under rotation about the $x$-axis, integrating over the azimuthal factor of solid angle gives
$\sigma_{t}\psi(\mu)=\frac{1}{2}\left(\frac{\sigma_{s,0}}{\gamma(1-\beta\mu)^{3}}\right)\int_{-1}^{1}(1-\beta\mu^{\prime})^{2}\psi(\mu^{\prime})d\mu^{\prime}\;\;.$ (9)
Equation (9) implies
$R_{s}(\mu^{\prime}\rightarrow\mu)=\frac{1}{2}\left(\frac{\sigma_{s,0}}{\gamma(1-\beta\mu)^{3}}\right)(1-\beta\mu^{\prime})^{2}$
in Eq. (1), which would become
$\frac{1}{c}\frac{\partial\psi}{\partial t}+\mu\frac{\partial\psi}{\partial x}+\sigma_{t}\psi=\frac{1}{2}\left(\frac{\sigma_{s,0}}{\gamma(1-\beta\mu)^{3}}\right)\int_{-1}^{1}(1-\beta\mu^{\prime})^{2}\psi(\mu^{\prime})d\mu^{\prime}+\frac{q}{2}\;\;,$ (10)
or
$\frac{1}{c}\frac{\partial\psi}{\partial t}+\mu\frac{\partial\psi}{\partial x}+\sigma_{t,0}\gamma(1-\beta\mu)\psi=\frac{1}{2}\left(\frac{\sigma_{s,0}}{\gamma(1-\beta\mu)^{3}}\right)\int_{-1}^{1}(1-\beta\mu^{\prime})^{2}\psi(\mu^{\prime})d\mu^{\prime}+\frac{q}{2}\;\;.$ (11)
For $\beta=0$, Eq. (11) reduces to the usual grey transport equation with isotropic scattering.
2.2 $\lambda$ function and $\Lambda$ integral
For convenience, we introduce the class of functions,
$\lambda_{n,k}(\mu)=\frac{\mu^{k}}{\gamma^{n}(1-\beta\mu)^{n}}\;\;.$ (12)
We find that using and reasoning about these functions in the context of the following sections and appendix expedites evaluation of relativistic factors. Here we review the properties of this class of functions through multiplication, differentiation in space or time, and integration over $\mu$; these operations are relevant to the following sections and appendix.
We note Eq. (12) readily satisfies a simple product rule that can be useful for keeping track of multiple Lorentz transforms in products of terms in harmonic ($\mu^{k}$-weighted) expansions,
$\lambda_{n+m,k+l}(\mu)=\lambda_{n,k}(\mu)\lambda_{m,l}(\mu)\;\;.$ (13)
Differentiation in space or time can be posed as a recursion,
$\frac{\partial\lambda_{n,k}}{\partial x}=\mu^{k}\left(\gamma^{-n}\frac{\partial(1-\beta\mu)^{-n}}{\partial x}+(1-\beta\mu)^{-n}\frac{\partial\gamma^{-n}}{\partial x}\right)=n\gamma\frac{\partial\beta}{\partial x}\left(\lambda_{n+1,k+1}(\mu)-\gamma\lambda_{n,k}(\mu)\right)\;\;,$ (14)
where we have made use of
$\frac{\partial\gamma}{\partial x}=\gamma^{3}\frac{\partial\beta}{\partial x}\;\;.$ (15)
In Section 2.4 we observe occurrences of
$\frac{\partial\ln\lambda_{n,k}}{\partial x}=\frac{1}{\lambda_{n,k}}\frac{\partial\lambda_{n,k}}{\partial x}=n\gamma\frac{\partial\beta}{\partial x}\left(\lambda_{1,1}(\mu)-\gamma\right)\;\;,$ (16)
which can be seen by applying Eq. (13). We caution that derivatives of $\lambda_{n,k}$ with respect to space or time are more complicated to evaluate in terms of $\mu_{0}$, given $\mu_{0}$ varies in space and time (see Section 2.3).
We further introduce integrals of $\lambda_{n,k}(\mu)$ over $\mu$,
$\Lambda_{n,k}=\int_{-1}^{1}\lambda_{n,k}(\mu)d\mu\;\;.$ (17)
It is straightforward to show (Appendix Section A) that the $\Lambda_{n,k}$ values follow a triangular recursion relationship,
$\Lambda_{n,k}=\frac{1}{\beta}\Lambda_{n,k-1}-\frac{1}{\gamma\beta}\Lambda_{n-1,k-1}\;\;.$ (18)
We may use this recursion to efficiently evaluate the 1D planar lab-frame harmonic Pn-series expansion of Eq. (11), as shown in the Appendix.
By virtue of the Lorentz transform for frequency, we also note the relation,
$\lambda_{n,0}(\mu)=\frac{1}{\gamma^{n}(1-\beta\mu)^{n}}=\gamma^{n}(1+\beta\mu_{0})^{n}=\frac{1}{\lambda_{n,0}(-\mu_{0})}\;\;,$ (19)
and for $k>0$,
$\lambda_{n,k}(\mu)=\frac{\mu^{k}}{\gamma^{n}(1-\beta\mu)^{n}}=\frac{(\mu_{0}+\beta)^{k}}{(1+\beta\mu_{0})^{k}}\gamma^{n}(1+\beta\mu_{0})^{n}=\gamma^{k}\sum_{q=0}^{k}\binom{k}{q}\frac{(-1)^{q}\beta^{k-q}}{\lambda_{n-k,-q}(-\mu_{0})}\;\;.$ (20)
We use Eqs. (19) and (20) in Section 2.4 to change reference frames before integration over $\mu_{0}$.
2.3 Angular derivative in the comoving frame
The approaches to asymptotic scaling makes use of the comoving frame. As such, it is worth noting that the space or time derivative of the comoving intensity, $\psi_{0}$, must account for the change in $\mu_{0}$ with space or time. Furthermore, from the Lorentz invariant relation of $\psi$ to $\psi_{0}$,
$\psi(x,\mu)=\frac{\psi_{0}(x,\mu_{0})}{\gamma^{4}(1-\mu\beta)^{4}}=\lambda_{4,0}(\mu)\psi_{0}(x,\mu_{0})\;\;.$ (21)
Using the angular derivative, the relationship between spatial derivatives is
$\frac{\partial\psi}{\partial x}=\frac{\partial}{\partial x}\left(\lambda_{4,0}\psi_{0}\right)=\psi_{0}\frac{\partial\lambda_{4,0}}{\partial x}+\lambda_{4,0}\left(\left.\frac{\partial\psi_{0}}{\partial x}\right|_{\mu_{0}}+\frac{\partial\mu_{0}}{\partial x}\frac{\partial\psi_{0}}{\partial\mu_{0}}\right)\;\;,$ (22)
where $|_{\mu_{0}}$ indicates $\mu_{0}$ is fixed when taking the partial derivative in $x$ (in order to avoid abuse of notation). The time derivative can similarly be found, resulting in $x$ replaced with $t$ in Eq. (22). In 1D planar geometry a simple equation for $\mu_{0}$ exists in terms of $\mu$ and $\beta$,
$\mu_{0}=\frac{\mu-\beta}{1-\beta\mu}\;\;.$ (23)
Consequently, assuming a given space-time dependence of $\beta$,
$\frac{\partial\mu_{0}}{\partial x}=-\frac{(1-\mu^{2})}{(1-\beta\mu)^{2}}\frac{\partial\beta}{\partial x}=-\gamma^{2}(1-\mu_{0}^{2})\frac{\partial\beta}{\partial x}\;\;,$ (24)
where we have used the fact that $\mu$ is constant in space and time in the lab frame, for a particular solution $\psi(x,\mu)$. The time derivative can again be similarly found, resulting in $x$ replaced with $t$ in Eq. (24).
2.4 Asymptotic scaling
Here we present two scaling methods: in the comoving frame using the standard approach for static background matter, and with a modification that includes an anisotropic scalar multiple of $\psi$. We first attempt to demonstrate the former does not furnish a closed equation to leading order, then we attempt to show the latter does give a closed drift-diffusion equation. In both attempts, we assume the existence of a “smallness” parameter $\varepsilon$ that is applicable to both the comoving and lab frame, and that intensity can be expanded in powers of $\varepsilon$,
$\psi=\sum_{k=0}^{\infty}\psi^{(k)}\varepsilon^{k}\;\;,$ (25a)
$\psi_{0}=\sum_{k=0}^{\infty}\psi_{0}^{(k)}\varepsilon^{k}\;\;.$ (25b)
We furthermore assume that Eq. (21) holds for each order,
$\psi^{(k)}(x,\mu)=\lambda_{4,0}(\mu)\psi_{0}^{(k)}(x,\mu_{0})\;\;,$ (26)
which recovers the full intensity invariance. We must impose Eq. (26) for $k\in\{0,1,2\}$ in order to undertake the transformations from the lab to the comoving frame in the following sections.
Concerning the $\varepsilon$-scaling of transport equation, in an equation with a lab-frame time derivative, we consider scaling of the form
$A[\psi]=\frac{1}{c}\frac{\partial\psi}{\partial t}=\frac{1}{c}\frac{\partial\psi}{\partial t}+\beta\frac{\partial\psi}{\partial x}+B(\mu)\psi-\beta\frac{\partial\psi}{\partial x}-B(\mu)\psi\\ \rightarrow\varepsilon\left(\frac{1}{c}\frac{\partial\psi}{\partial t}+\beta\frac{\partial\psi}{\partial x}+B(\mu)\psi\right)-\beta\frac{\partial\psi}{\partial x}-B(\mu)\psi=A_{\varepsilon}[\psi]\;\;,$ (27)
where $A[\cdot]$ is the original time-derivative operator (or the remainder of the transport equation), $B(\nu)$ is an as-yet unknown coefficient and $A_{\varepsilon}[\cdot]$ is a $\varepsilon$-scaled form of this operator. This scaling assumes that the residual excluded from the $\varepsilon$ coefficient scales as a leading-order contribution. The full set of scaling relations we use is
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}+B(\mu)\right)\psi\rightarrow\varepsilon\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}+B(\mu)\right)\psi\;\;,$ (28a)
$\sigma_{t,0}\rightarrow\frac{\sigma_{t,0}}{\varepsilon}\;\;,$ (28b)
$\sigma_{a,0}\rightarrow\varepsilon\sigma_{a,0}\;\;,$ (28c)
$q\rightarrow\varepsilon q\;\;.$ (28d)
In the following sections, we determine the form of $B(\mu)$ such that the equation for the leading-order scalar intensity is closed at O($\varepsilon^{2}$).
2.4.1 Scaling the comoving time derivative
Assuming $B(\mu)=0$, Eqs. (28) reduce to the standard asymptotic scaling for linear transport in the diffusion limit, in the comoving frame. Applying this to Eq. (1), and substituting in $\lambda_{n,0}$-functions,
$\varepsilon^{2}\left(\frac{1}{c}\frac{\partial\psi}{\partial t}+\beta\frac{\partial\psi}{\partial x}\right)+\varepsilon(\mu-\beta)\frac{\partial\psi}{\partial x}+\frac{\sigma_{t,0}}{\lambda_{1,0}}\psi=\frac{1}{2}\lambda_{3,0}(\sigma_{t,0}-\varepsilon^{2}\sigma_{a,0})\int_{-1}^{1}\frac{\psi(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}+\varepsilon^{2}\frac{q}{2}\;\;,$ (29)
where $\lambda_{n,0}=\lambda_{n,0}(\mu)$ for brevity. Incorporating Eq. (25) into Eq. (29) and matching coefficients of $\varepsilon^{k}$, the O(1) equation is
$\psi^{(0)}=\frac{1}{2}\lambda_{4,0}\int_{-1}^{1}\frac{\psi^{(0)}(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}\;\;,$ (30)
and the O($\varepsilon$) equation is
$\frac{(\mu-\beta)}{(1-\beta\mu)}\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\psi^{(0)}}{\partial x}+\psi^{(1)}=\frac{1}{2}\lambda_{4,0}\int_{-1}^{1}\frac{\psi^{(1)}(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}\;\;.$ (31)
The O(1) equation is a statement of isotropy in the comoving frame, while the O($\varepsilon$) reduces to the usual linearly anisotropic form when $\beta=0$.
Converting $\mu$, $\psi^{(0)}$ and $\psi^{(1)}$ to the comoving frame and dividing by a factor of $\lambda_{4,0}(\mu)$, Eqs. (30) and (31) become
$\psi_{0}^{(0)}=\frac{1}{2}\phi_{0}^{(0)}\;\;,$ (32)
and
$\frac{1}{\gamma\sigma_{t,0}}\mu_{0}\left(\frac{\psi_{0}^{(0)}}{\lambda_{4,0}}\frac{\partial\lambda_{4,0}}{\partial x}+\left.\frac{\partial\psi_{0}^{(0)}}{\partial x}\right|_{\mu_{0}}-\gamma^{2}\frac{\partial\beta}{\partial x}(1-\mu_{0}^{2})\frac{\partial\psi_{0}^{(0)}}{\partial\mu_{0}}\right)+\psi_{0}^{(1)}=\frac{1}{2}\phi_{0}^{(1)}\;\;.$ (33)
Given $\psi_{0}^{(0)}$ is isotropic, the angular derivative vanishes and $|_{\mu_{0}}$ can be removed from the derivative, leaving
$\frac{1}{\gamma\sigma_{t,0}}\mu_{0}\left(\psi_{0}^{(0)}\frac{\partial\ln\lambda_{4,0}}{\partial x}+\frac{\partial\psi_{0}^{(0)}}{\partial x}\right)+\psi_{0}^{(1)}=\frac{1}{2}\phi_{0}^{(1)}\;\;.$ (34)
If $\beta$ is constant in space, Eq. (34) reduces to the usual linear anisotropic form in the comoving frame (consistent with the lab-frame form of the equation). If $\beta$ is not constant in space, integrating over $\mu_{0}$ and simplifying gives
$\frac{4}{3}\gamma^{2}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}=0\;\;,$ (35)
which is a contradiction, unless $\phi_{0}^{(0)}=0$.
2.4.2 Scaling the comoving time derivative with the anisotropic term
The contradiction resulting from scaling the comoving time derivative suggests that we may find a non-zero form of $B(\mu)$ that eliminates the anisotropic term proportional to the velocity gradient. Here we show that a form that accomplishes this is
$B(\mu)=(\mu-\beta)\frac{\partial}{\partial x}\ln(\lambda_{4,0}(\mu))\;\;.$ (36)
Applying the scaling with Eq. (36),
$\varepsilon^{2}\left(\frac{1}{c}\frac{\partial\psi}{\partial t}+\beta\frac{\partial\psi}{\partial x}+\psi\,(\mu-\beta)\frac{\partial\ln\lambda_{4,0}}{\partial x}\right)\\ +\varepsilon(\mu-\beta)\left(\frac{\partial\psi}{\partial x}-\psi\,\frac{\partial\ln\lambda_{4,0}}{\partial x}\right)+\frac{\sigma_{t,0}}{\lambda_{1,0}}\psi=\\ \frac{1}{2}\lambda_{3,0}(\sigma_{t,0}-\varepsilon^{2}\sigma_{a,0})\int_{-1}^{1}\frac{\psi(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}+\varepsilon^{2}\frac{q}{2}\;\;.$ (37)
Incorporating Eq. (25) into Eq. (37) and matching coefficients of $\varepsilon^{k}$, the O(1) equation is
$\psi^{(0)}=\frac{1}{2}\lambda_{4,0}\int_{-1}^{1}\frac{\psi^{(0)}(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}\;\;,$ (38)
and the O($\varepsilon$) equation is
$\frac{(\mu-\beta)}{(1-\beta\mu)}\frac{1}{\gamma\sigma_{t,0}}\left(\frac{\partial\psi^{(0)}}{\partial x}-\psi^{(0)}\,\frac{\partial\ln\lambda_{4,0}}{\partial x}\right)+\psi^{(1)}=\frac{1}{2}\lambda_{4,0}\int_{-1}^{1}\frac{\psi^{(1)}(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}\;\;.$ (39)
Converting $\mu$, $\psi^{(0)}$ and $\psi^{(1)}$ to the comoving frame and dividing by a factor of $\lambda_{4,0}(\mu)$, Eqs. (38) and (39) become
$\psi_{0}^{(0)}=\frac{1}{2}\phi_{0}^{(0)}\;\;,$ (40)
and
$\frac{1}{\gamma\sigma_{t,0}}\mu_{0}\left(\frac{\psi_{0}^{(0)}}{\lambda_{4,0}}\frac{\partial\lambda_{4,0}}{\partial x}+\frac{\partial\psi_{0}^{(0)}}{\partial x}-\psi_{0}^{(0)}\,\frac{\partial\ln\lambda_{4,0}}{\partial x}\right)+\psi_{0}^{(1)}=\frac{1}{2}\phi_{0}^{(1)}\;\;,$ (41)
where we have made use again of isotropy to eliminate the angular derivative. Simplifying the O($\varepsilon$) equation gives
$\frac{1}{\gamma\sigma_{t,0}}\mu_{0}\frac{\partial\psi_{0}^{(0)}}{\partial x}+\psi_{0}^{(1)}=\frac{1}{2}\phi_{0}^{(1)}\;\;,$ (42)
which is now the standard linear anisotropic relationship, posed in terms of comoving quantities. The contradiction encountered by scaling the comoving time derivative is no longer manifest in Eq. (42).
Proceeding to the O($\varepsilon^{2}$) equation, matching $\varepsilon^{2}$ coefficients gives
$\lambda_{1,0}\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\psi^{(0)}+\psi^{(0)}\frac{(\mu-\beta)}{\gamma(1-\beta\mu)}\frac{\partial\ln\lambda_{4,0}}{\partial x}\\ +\frac{(\mu-\beta)}{\gamma(1-\beta\mu)}\left(\frac{\partial\psi^{(1)}}{\partial x}-\psi^{(1)}\,\frac{\partial\ln\lambda_{4,0}}{\partial x}\right)+\sigma_{t,0}\psi^{(2)}=\\ \frac{1}{2}\lambda_{4,0}\left(\sigma_{t,0}\int_{-1}^{1}\frac{\psi^{(2)}(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}-\sigma_{a,0}\int_{-1}^{1}\frac{\psi^{(0)}(\mu^{\prime})}{\lambda_{2,0}(\mu^{\prime})}d\mu^{\prime}\right)+\lambda_{1,0}\frac{q}{2}\;\;.$ (43)
Converting this to the comoving frame and dividing by $\lambda_{4,0}$,
$\lambda_{1,0}\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\psi_{0}^{(0)}+\lambda_{1,0}\psi_{0}^{(0)}\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\ln\lambda_{4,0}+\psi_{0}^{(0)}\frac{\mu_{0}}{\gamma}\frac{\partial\ln\lambda_{4,0}}{\partial x}\\ +\frac{\mu_{0}}{\gamma}\left(\left.\frac{\partial\psi_{0}^{(1)}}{\partial x}\right|_{\mu_{0}}-\gamma^{2}(1-\mu_{0}^{2})\frac{\partial\beta}{\partial x}\frac{\partial\psi_{0}^{(1)}}{\partial\mu_{0}}+\frac{\psi_{0}^{(1)}}{\lambda_{4,0}}\frac{\partial\lambda_{4,0}}{\partial x}-\psi_{0}^{(1)}\,\frac{\partial\ln\lambda_{4,0}}{\partial x}\right)+\sigma_{t,0}\psi_{0}^{(2)}=\\ \frac{1}{2}\left(\sigma_{t,0}\phi_{0}^{(2)}-\sigma_{a,0}\phi_{0}^{(0)}\right)+\frac{q}{2\lambda_{3,0}}\;\;,$ (44)
where use has been made of $\psi_{0}^{(0)}$ isotropy in evaluating the first term on the left side, to eliminate derivaties with respect to $\mu_{0}$. We may simplify the $k=1$ terms by canceling the $B(\mu)$ residue and using Eq. (42),
$\lambda_{1,0}\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\psi_{0}^{(0)}+\lambda_{1,0}\psi_{0}^{(0)}\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\ln\lambda_{4,0}+\psi_{0}^{(0)}\frac{\mu_{0}}{\gamma}\frac{\partial\ln\lambda_{4,0}}{\partial x}\\ +\frac{\mu_{0}}{\gamma}\frac{\partial}{\partial x}\left(\frac{\phi_{0}^{(1)}}{2}\right)-\frac{\mu_{0}}{\gamma}\left.\frac{\partial}{\partial x}\left(\frac{\mu_{0}}{\gamma\sigma_{t,0}}\frac{\partial\psi_{0}^{(0)}}{\partial x}\right)\right|_{\mu_{0}}+\mu_{0}(1-\mu_{0}^{2})\frac{\partial\beta}{\partial x}\frac{1}{\sigma_{t,0}}\frac{\partial\psi_{0}^{(0)}}{\partial x}+\sigma_{t,0}\psi_{0}^{(2)}=\\ \frac{1}{2}\left(\sigma_{t,0}\phi_{0}^{(2)}-\sigma_{a,0}\phi_{0}^{(0)}\right)+\frac{q}{2\lambda_{3,0}}\;\;.$ (45)
Applying Eq. (16) to the derivatives of $\ln\lambda_{4,0}$,
$\lambda_{1,0}\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\psi_{0}^{(0)}+4\left[\gamma\left(\lambda_{2,1}-\gamma\lambda_{1,0}\right)\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta+\mu_{0}\left(\lambda_{1,1}-\gamma\right)\frac{\partial\beta}{\partial x}\right]\psi_{0}^{(0)}\\ +\frac{\mu_{0}}{\gamma}\frac{\partial}{\partial x}\left(\frac{\phi_{0}^{(1)}}{2}\right)-\frac{\mu_{0}^{2}}{\gamma}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\psi_{0}^{(0)}}{\partial x}\right)+\mu_{0}(1-\mu_{0}^{2})\frac{\partial\beta}{\partial x}\frac{1}{\sigma_{t,0}}\frac{\partial\psi_{0}^{(0)}}{\partial x}+\sigma_{t,0}\psi_{0}^{(2)}=\\ \frac{1}{2}\left(\sigma_{t,0}\phi_{0}^{(2)}-\sigma_{a,0}\phi_{0}^{(0)}\right)+\frac{q}{2\lambda_{3,0}}\;\;.$ (46)
Now no occurrences of $\lambda$-functions and $\mu_{0}$ are inside space or time derivatives. The source term $q$ may be related to the comoving source term by
$q(\mu)=q_{0}(\mu_{0})\lambda_{3,0}(\mu)\;\;,$
where $\lambda_{3,0}$ arises from lab-frame frequency integration of the Lorentz invariant relationship for emissivity. Substituting $q_{0}$ for $q$ and using Eq. (20) we may convert the $\lambda$-functions to polynomials in $\mu_{0}$,
$\gamma(1+\beta\mu_{0})\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\psi_{0}^{(0)}+4\gamma\left(\mu_{0}+\beta-1\right)\left[\gamma^{2}(1+\beta\mu_{0})\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta+\mu_{0}\frac{\partial\beta}{\partial x}\right]\psi_{0}^{(0)}\\ +\frac{\mu_{0}}{\gamma}\frac{\partial}{\partial x}\left(\frac{\phi_{0}^{(1)}}{2}\right)-\frac{\mu_{0}^{2}}{\gamma}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\psi_{0}^{(0)}}{\partial x}\right)+\mu_{0}(1-\mu_{0}^{2})\frac{\partial\beta}{\partial x}\frac{1}{\sigma_{t,0}}\frac{\partial\psi_{0}^{(0)}}{\partial x}+\sigma_{t,0}\psi_{0}^{(2)}=\\ \frac{1}{2}\left(\sigma_{t,0}\phi_{0}^{(2)}-\sigma_{a,0}\phi_{0}^{(0)}\right)+\frac{q_{0}}{2}\;\;.$ (47)
Integrating Eq. (47) over $\mu_{0}$, dividing by a factor of $\gamma$ and simplifying gives
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+4\left[\gamma^{2}\left(\frac{4}{3}\beta-1\right)\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta+\frac{1}{3}\frac{\partial\beta}{\partial x}\right]\phi_{0}^{(0)}\\ -\frac{1}{3\gamma^{2}}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\frac{\sigma_{a,0}}{\gamma}\phi_{0}^{(0)}+\frac{1}{2\gamma}\int_{-1}^{1}q_{0}(\mu_{0})d\mu_{0}\;\;.$ (48)
We see that we have obtained a drift-diffusion equation for the leading-order scalar intensity using our lab-frame asymptotic scaling with $B(\mu)$ from Eq. (36). Furthermore, we note that we have not assumed a constraint on the value of $\beta$ itself. The diffusion stencil has three factors of $1/\gamma$, consistent with the finding of [giona2017] for bi-directional 1D planar random walk. However, we have an additional $\beta$-dependent coefficient of $\phi_{0}^{(0)}$, which accounts for the energetic effect of Doppler shift and adiabatic expansion. To more clearly see this, we may neglect the time-dependence of $\beta$, and neglect terms of O($\beta^{2}$) and O($\beta\partial\beta/\partial x$), to obtain
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3}\frac{\partial}{\partial x}\left(\frac{1}{\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\sigma_{a,0}\phi_{0}^{(0)}+\frac{1}{2}\int_{-1}^{1}q_{0}(\mu_{0})d\mu_{0}\;\;,$ (49)
which is the standard semi-relativistic form (see, for instance, [castor2004], Chapter 6). If $\beta$ is constant and approaches 1, assuming $q$ and $\phi_{0}^{(0)}$ remain bounded,
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}=0\;\;,$ (50)
which is the limiting behavior observed by [giona2017].
2.4.3 Scaling the Lagrangian time derivative of $\beta$
We note that there is pathology in Eq. (48): the coefficient accounting for adiabatic and Doppler shift has a factor that is mixed-parity in $\beta$, $(4\beta/3-1)$, multiplied to the Lagrangian derivative of $\beta$. This combination of factors breaks symmetry under the parity transformation $(x,\beta)\rightarrow(-x,-\beta)$, which is maintained in the semi-relativistic form of Eq. (49). To preserve this physical symmetry, this suggests we must impose an asymptotic non-acceleration condition for the matter, for instance,
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta\rightarrow\varepsilon\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta\;\;.$ (51)
Incorporating Eq. (51) into the suite of scaling conditions removes the term corresponding to the Lagrangian derivative of velocity in the O($\varepsilon^{2}$) equation, resulting in
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3\gamma^{2}}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\frac{\sigma_{a,0}}{\gamma}\phi_{0}^{(0)}+\frac{Q_{0}}{\gamma}\;\;.$ (52)
where we have replaced the integral of the comoving external source with $Q_{0}$. This equation now has the same coefficient for adiabatic and expansion effects as the standard semi-relativistic form, so retains mirrored behavior in scalar intensity when $(x,\beta)\rightarrow(-x,-\beta)$. In subsequent sections we examine Eq. (52), which corresponds to the following list of lab-frame scaling relations, now including Eq. (51),
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}+(\mu-\beta)\frac{\partial}{\partial x}\ln(\lambda_{4,0}(\mu))\right)\psi\rightarrow\varepsilon\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}+(\mu-\beta)\frac{\partial}{\partial x}\ln(\lambda_{4,0}(\mu))\right)\psi\;\;,$ (53a)
$\sigma_{t,0}\rightarrow\frac{\sigma_{t,0}}{\varepsilon}\;\;,$ (53b)
$\sigma_{a,0}\rightarrow\varepsilon\sigma_{a,0}\;\;,$ (53c)
$q\rightarrow\varepsilon q\;\;,$ (53d)
$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta\rightarrow\varepsilon\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\beta\;\;.$ (53e)
3 Discretization
3.1 Spatial stencil
Here we describe the discretization of Eq. (48) we use to explore properties of the solutions in Section 4. We solve the resulting discretized system deterministically, but we note that there exists a straightforward Discrete Diffusion Monte Carlo (DDMC) [gentile2001, densmore2007, abdikamalov2012] interpretation of the discretized terms, and write them following the notation of [densmore2007]. In 1D, a grey deterministic solution is straightforward to implement, and we do not seek to hybridize lab-frame MC transport with DDMC in the scope of this work.
We make the assumption that $\beta$ and the spatial derivatives of $\beta$ are evaluated per spatial cell, prior to spatial discretization of Eq. (48). Similar to opacity, we assume $\beta$ and its derivatives are given with a time step and spatial cell. To simplify the presentation of the spatial discretization that follows, we introduce a label for the coefficient of $\phi_{0}^{(0)}$ on the left side,
$\eta_{i}=\frac{4}{3}\left.\frac{\partial\beta}{\partial x}\right|_{i}$ (54)
where subscript $i$ is the index of the spatial cell, indicating evaluation at the cell.
Finite-volume discretization of Eq. (48) gives
$\frac{1}{c}\frac{\partial\phi_{0,i}}{\partial t}+\frac{1}{\Delta x_{i}}\left(\beta_{i+1/2}\phi_{0,i+1/2}-\beta_{i-1/2}\phi_{0,i-1/2}\right)+\eta_{i}\phi_{0,i}+\frac{1}{\gamma_{i}}\sigma_{a,0,i}\phi_{0,i}\\ -\frac{1}{\gamma_{i}^{2}\Delta x_{i}}\left(\frac{1}{3\gamma_{i+1/2}\sigma_{t,0,i+1/2}}\left.\frac{\partial\phi_{0}}{\partial x}\right|_{i+1/2}-\frac{1}{3\gamma_{i-1/2}\sigma_{t,0,i-1/2}}\left.\frac{\partial\phi_{0}}{\partial x}\right|_{i-1/2}\right)=\frac{Q_{0,i}}{\gamma_{i}}\;\;,$ (55)
where we have dropped the superscript $(0)$, understanding that our leading-order solution will be the full solution for the diffusion equation. Consequently, we need to find $\phi_{0,i\pm 1/2}$ in terms of $\phi_{0,i}$ and $\phi_{0,i\pm 1}$ from auxiliary equations.
A simple way to treat the advection portion of the Lagrangian time-derivative is to upwind $\phi_{0,i\pm 1/2}$ based on the sign of $\beta_{i\pm 1/2}$,
$\phi_{0,i+1/2}=\Theta(\beta_{i+1/2})\phi_{0,i}+(1-\Theta(\beta_{i+1/2}))\phi_{0,i+1}\;\;,$ (56)
where $\Theta(\cdot)$ is the standard unit step (or Heaviside) function. Equation (56) acts either like a sink ($\beta_{i+1/2}>0$) or a source ($\beta_{i+1/2}<0$), but not both. If we permit discontinuous velocity at cell edges, introducing superscript $\pm$ to denote evaluation immediately below (-) or above (+) the edge along $x$, we may generalize Eq. (56),
$\beta_{i+1/2}\phi_{0,i+1/2}=\Theta(\beta_{i+1/2}^{-})\beta_{i+1/2}^{-}\phi_{0,i}+\Theta(-\beta_{i+1/2}^{+})\beta_{i+1/2}^{+}\phi_{0,i+1}\;\;,$ (57a)
$\beta_{i-1/2}\phi_{0,i-1/2}=\Theta(-\beta_{i-1/2}^{+})\beta_{i-1/2}^{+}\phi_{0,i}+\Theta(\beta_{i-1/2}^{-})\beta_{i-1/2}^{-}\phi_{0,i-1}\;\;,$ (57b)
which permits simultaneous sources and sinks at the cell edge $i+1/2$, and reduces to Eq. (56) when the velocity is continuous ($\beta_{i+1/2}^{-}=\beta_{i+1/2}^{+}=\beta_{i+1/2}$). Incorporating Eq. (57) into the advection term in Eq. (55),
$\frac{1}{\Delta x_{i}}\left(\beta_{i+1/2}\phi_{0,i+1/2}-\beta_{i-1/2}\phi_{0,i-1/2}\right)=\\ \frac{1}{\Delta x_{i}}\left(\Theta(\beta_{i+1/2}^{-})\beta_{i+1/2}^{-}\phi_{0,i}-\Theta(-\beta_{i+1/2}^{+})|\beta_{i+1/2}^{+}|\phi_{0,i+1}+\Theta(-\beta_{i-1/2}^{+})|\beta_{i-1/2}^{+}|\phi_{0,i}-\Theta(\beta_{i-1/2}^{-})\beta_{i-1/2}^{-}\phi_{0,i-1}\right)\\ =(\sigma_{A,i\rightarrow i+1}+\sigma_{A,i\rightarrow i-1})\phi_{0,i}-\frac{\Delta x_{i+1}}{\Delta x_{i}}\sigma_{A,i+1\rightarrow i}\phi_{0,i+1}-\frac{\Delta x_{i-1}}{\Delta x_{i}}\sigma_{A,i-1\rightarrow i}\phi_{0,i-1}\;\;,$ (58)
where
$\sigma_{A,i\rightarrow i-1}=\frac{1}{\Delta x_{i}}\Theta(-\beta_{i-1/2}^{+})|\beta_{i-1/2}^{+}|\;\;,$ (59a)
$\sigma_{A,i\rightarrow i+1}=\frac{1}{\Delta x_{i}}\Theta(\beta_{i+1/2}^{-})\beta_{i+1/2}^{-}\;\;,$ (59b)
can be viewed as advection “leakage opacities” [densmore2007] (see [abdikamalov2012] for an operator-split continuous treatment of advection).
The diffusion operator terms can be evaluated in the usual way in the domain interior [densmore2007],
$\frac{1}{3\gamma_{i+1/2}\sigma_{t,0,i+1/2}}\left.\frac{\partial\phi_{0}}{\partial x}\right|_{i+1/2}=\frac{2(\phi_{0,i+1}-\phi_{0,i})}{3(\gamma_{i+1/2}^{-}\sigma_{t,0,i+1/2}^{-}\Delta x_{i}+\gamma_{i+1/2}^{+}\sigma_{t,0,i+1/2}^{+}\Delta x_{i+1})}\;\;,$ (60)
where $\pm$-superscripts again indicate evaluation just to the lower or upper side of the cell edge $i+1/2$. The formulation permits several options for discretizing the velocity field, for instance: if $\gamma$ is continuous then $\gamma_{i+1/2}^{-}=\gamma_{i+1/2}^{+}=\gamma_{i+1/2}$; if $\gamma$ is piecewise-constant, then one may set $\gamma_{i+1/2}^{-}=\gamma_{i}$ and $\gamma_{i+1/2}^{+}=\gamma_{i+1}$. Consequently, the diffusion leakage opacities, corresponding to the standard non-relativistic form [densmore2007], are
$\sigma_{D,i\rightarrow i-1}=\frac{2}{3\gamma_{i}^{2}\Delta x_{i}}\left(\frac{1}{\gamma_{i-1/2}^{+}\sigma_{t,0,i-1/2}^{+}\Delta x_{i}+\gamma_{i-1/2}^{-}\sigma_{t,0,i-1/2}^{-}\Delta x_{i-1}}\right)\;\;,$ (61a)
$\sigma_{D,i\rightarrow i+1}=\frac{2}{3\gamma_{i}^{2}\Delta x_{i}}\left(\frac{1}{\gamma_{i+1/2}^{-}\sigma_{t,0,i+1/2}^{-}\Delta x_{i}+\gamma_{i+1/2}^{+}\sigma_{t,0,i+1/2}^{+}\Delta x_{i+1}}\right)\;\;.$ (61b)
Incorporating Eqs. (59) and (61) into Eq. (55),
$\frac{1}{c}\frac{\partial\phi_{0,i}}{\partial t}+\left(\sigma_{A,i\rightarrow i+1}+\sigma_{A,i\rightarrow i-1}+\sigma_{D,i\rightarrow i+1}+\sigma_{D,i\rightarrow i-1}+\eta_{i}+\frac{1}{\gamma_{i}}\sigma_{a,0,i}\right)\phi_{0,i}\\ =\frac{Q_{0,i}}{\gamma_{i}}+\frac{\Delta x_{i+1}}{\Delta x_{i}}\left(\sigma_{A,i+1\rightarrow i}+\frac{\gamma_{i+1}^{2}}{\gamma_{i}^{2}}\sigma_{D,i+1\rightarrow i}\right)\phi_{0,i+1}+\frac{\Delta x_{i-1}}{\Delta x_{i}}\left(\sigma_{A,i-1\rightarrow i}+\frac{\gamma_{i-1}^{2}}{\gamma_{i}^{2}}\sigma_{D,i-1\rightarrow i}\right)\phi_{0,i-1}\;\;.$ (62)
Equation (62) is very similar to a standard grey diffusion stencil, but with the addition of advection leakage, the $\eta$ coefficient, and $\gamma$ factors multiplying diffusion leakage opacities.
Finally, to treat the boundary, we assume the asymptotic boundary layer condition used by [densmore2007] holds in the comoving frame,
$2\int_{0}^{1}W(\mu_{0})\psi_{0}^{-}(x_{i-1/2},\mu_{0})d\mu_{0}=\phi_{0}(x_{i-1/2})-\frac{\lambda}{\gamma\sigma_{t,0,i-1/2}}\left.\frac{\partial\phi_{0}}{\partial x}\right|_{i-1/2}\;\;,$ (63)
where $W(\cdot)$ is the angular weighting function for intensity incident just below $x_{i-1/2}$, $\psi_{0}^{-}$, and $\lambda$ is the extrapolation distance [habetler1975]. Equation (63) is consistent with applying scaling Eqs. (47) to the transport equation, keeping terms up to O($\varepsilon$), transforming intensity and opacity to the comoving frame, and then applying the boundary layer analysis of [habetler1975] (the factor of $\gamma$ from the transformation of $\sigma_{t}$ to $\sigma_{t,0}$ can be included in the transformation of the spatial coordinate to the stretched optical depth parameter). Equation (63) can also be obtained from the the static boundary condition specified in comoving space-time, assuming the comoving time derivative is 0. Following the discretization procedure of [densmore2007], the leakage opacity and boundary transmission probability are
$\tilde{\sigma}_{D,i\rightarrow i-1}=\frac{2}{3\gamma_{i}^{2}\Delta x_{i}(\gamma_{i-1/2}^{+}\sigma_{t,0,i-1/2}^{+}\Delta x_{i}+2\lambda)}\;\;,$ (64a)
$P_{i-1/2}(\mu_{0})=\frac{4}{3\gamma_{i}^{2}(\gamma_{i-1/2}^{+}\sigma_{t,0,i-1/2}^{+}\Delta x_{i}+2\lambda)}\left(1+\frac{3}{2}\mu_{0}\right)\;\;,$ (64b)
which provide the boundary condition stencil for the full equation,
$\frac{1}{c}\frac{\partial\phi_{0,i}}{\partial t}+\left(\sigma_{A,i\rightarrow i+1}+\sigma_{A,i\rightarrow i-1}+\sigma_{D,i\rightarrow i+1}+\tilde{\sigma}_{D,i\rightarrow i-1}+\eta_{i}+\frac{1}{\gamma_{i}}\sigma_{a,0,i}\right)\phi_{0,i}\\ =\frac{Q_{0,i}}{\gamma_{i}}+\frac{\Delta x_{i+1}}{\Delta x_{i}}\left(\sigma_{A,i+1\rightarrow i}+\frac{\gamma_{i+1}^{2}}{\gamma_{i}^{2}}\sigma_{D,i+1\rightarrow i}\right)\phi_{0,i+1}\\ +\frac{1}{\Delta x_{i}}\left(\beta_{i-1/2}^{+}\phi_{0,i-1/2}^{-}+\int_{0}^{1}\mu_{0}P_{i-1/2}(\mu_{0})\psi_{0}^{-}(x_{i-1/2},\mu_{0})d\mu_{0}\right)\;\;.$ (65)
In a hybrid transport-diffusion scheme, a transporting MC particle incident at i1/2i-1/2 on the DDMC region can have its energy and direction transformed into the comoving frame, then admission into the DDMC region can be determined by $P_{i-1/2}(\mu_{0})$. We note that the $\gamma$-factors in the denominator act only to lower the probability relative to the non-relativistic form, and hence do not complicate the restriction $P_{i-1/2}(\mu_{0})\leq 1$.
3.2 Time stencil and stability
Considering the spatial stencil for the domain interior, and further assuming velocity and opacity are constant in a time step, a simple time discretization is
$\frac{\phi_{0,i,n+1}-\phi_{0,i,n}}{c\Delta t_{n}}+\left(\sigma_{A,i\rightarrow i+1}+\sigma_{A,i\rightarrow i-1}+\sigma_{D,i\rightarrow i+1}+\sigma_{D,i\rightarrow i-1}+\eta_{i}+\frac{1}{\gamma_{i}}\sigma_{a,0,i}\right)\bar{\phi}_{0,i}\\ =\frac{\bar{Q}_{0,i}}{\gamma_{i}}+\frac{\Delta x_{i+1}}{\Delta x_{i}}\left(\sigma_{A,i+1\rightarrow i}+\frac{\gamma_{i+1}^{2}}{\gamma_{i}^{2}}\sigma_{D,i+1\rightarrow i}\right)\bar{\phi}_{0,i+1}+\frac{\Delta x_{i-1}}{\Delta x_{i}}\left(\sigma_{A,i-1\rightarrow i}+\frac{\gamma_{i-1}^{2}}{\gamma_{i}^{2}}\sigma_{D,i-1\rightarrow i}\right)\bar{\phi}_{0,i-1}\;\;,$ (66)
where
$\bar{\phi}_{0,i}=(1-\alpha)\phi_{0,i,n}+\alpha\phi_{0,i,n+1}\;\;,$ (67)
and $\alpha\in[0,1]$ is a time-centering parameter (consequently, $\alpha=0,1/2,1$ correspond to explicit Euler, Crank-Nicolson, and implicit Euler schemes, respectively). We may rewrite Eq. (66) as
$\phi_{0,i,n+1}+\mathbf{A}_{i,i-1}\bar{\phi}_{0,i-1}+\mathbf{A}_{i,i}\bar{\phi}_{0,i}+\mathbf{A}_{i,i+1}\bar{\phi}_{0,i+1}=\phi_{0,i,n}+c\Delta t_{n}\frac{\bar{Q}_{0,i}}{\gamma_{i}}\;\;,$ (68)
where
$\mathbf{A}_{i,i}=c\Delta t_{n}\left(\sigma_{A,i\rightarrow i+1}+\sigma_{A,i\rightarrow i-1}+\sigma_{D,i\rightarrow i+1}+\sigma_{D,i\rightarrow i-1}+\eta_{i}+\frac{1}{\gamma_{i}}\sigma_{a,0,i}\right)\;\;,$ (69a)
$\mathbf{A}_{i,i\pm 1}=-\frac{\Delta x_{i\pm 1}}{\Delta x_{i}}c\Delta t_{n}\left(\sigma_{A,i\pm 1\rightarrow i}+\frac{\gamma_{i\pm 1}^{2}}{\gamma_{i}^{2}}\sigma_{D,i\pm 1\rightarrow i}\right)\;\;,$ (69b)
are entries of a tri-diagonal matrix, which due to $\eta_{i}$ may not be positive-definite. For uniform spatial cells, the $\mathbf{A}$-matrix is symmetric if $\beta_{i+1/2}^{-}=\beta_{i+1/2}^{+}$. Substituting in the right side of Eq. (67) for $\bar{\phi}_{0}$,
$\alpha\mathbf{A}_{i,i-1}\phi_{0,i-1,n+1}+\left(1+\alpha\mathbf{A}_{i,i}\right)\phi_{0,i,n+1}+\alpha\mathbf{A}_{i,i+1}\phi_{0,i+1,n}\\ =c\Delta t_{n}\frac{\bar{Q}_{0,i}}{\gamma_{i}}-(1-\alpha)\mathbf{A}_{i,i-1}\phi_{0,i-1,n}+\left(1-(1-\alpha)\mathbf{A}_{i,i}\right)\phi_{0,i,n}-(1-\alpha)\mathbf{A}_{i,i+1}\phi_{0,i+1,n}\;\;.$ (70)
We may perform a Fourier error analysis for Eq. (70). For a spatial wavenumber $K$, assuming $\phi_{0,i,n}^{(e)}$ is an exact solution to the stencil, substituting
$\phi_{0,i,n}=\phi_{0,i,n}^{(e)}+\delta\phi_{K,n}e^{iKx_{i}}\;\;,$ (71)
into Eq. (70) gives
$\frac{\delta\phi_{K,n+1}}{\delta\phi_{K,n}}=\frac{-(1-\alpha)\mathbf{A}_{i,i-1}e^{-iK\Delta x_{i-1/2}}+1-(1-\alpha)\mathbf{A}_{i,i}-(1-\alpha)\mathbf{A}_{i,i+1}e^{iK\Delta x_{i+1/2}}}{\alpha\mathbf{A}_{i,i-1}e^{-iK\Delta x_{i-1/2}}+1+\alpha\mathbf{A}_{i,i}+\alpha\mathbf{A}_{i,i+1}e^{iK\Delta x_{i+1/2}}}\;\;.$ (72)
We note we have abused notation by using $i$ as an imaginary number when multiplying $K$, and as a cell index subscript. The stability condition is that the modulus of Eq. (72) be less than or equal to 1. In Section 4, we only use a uniform spatial grid; setting $\theta=K\Delta x=K\Delta x_{i\pm 1/2}$ in Eq. (72),
$\frac{\delta\phi_{K,n+1}}{\delta\phi_{K,n}}=\frac{\mathcal{R}(\phi,\alpha-1)}{\mathcal{R}(\phi,\alpha)}\;\;,$ (73)
where
$\mathcal{R}(\theta,z)=1+z\mathbf{A}_{i,i}+z\mathbf{A}_{i,i-1}e^{-i\theta}+z\mathbf{A}_{i,i+1}e^{i\theta}\;\;.$ (74)
Figure 1 has example stability curves for $\beta=0.6$ and $\alpha=0$ (explicit Euler) or $\alpha=0.5$ (Crank-Nicolson), with $\Delta x,c\Delta t,\sigma_{t},\sigma_{a}=1/128,1/64,128,1/2$, respectively, and $\mathcal{G}(\theta,\alpha)=|\mathcal{R}(\phi,\alpha-1)/\mathcal{R}(\phi,\alpha)|$. Comparing to the static diffusion equation ($\beta=0$), we see the effect of time centering on stability depends on velocity: for $\alpha=0$ there is a significantly larger range of spatial wavenumbers that correspond to unstable error modes. The most significant contributor to the enhancement in instability for $\alpha=0$ are the advective terms resulting from the Lagrangian derivative, consistent with advection imposing a CFL-type condition on time step size.
Figure 1: Polar graphs of stability measure $\mathcal{G}(\theta,\alpha)$ for $\beta=0.6$ versus angle $\theta=K\Delta x$ (green dotted). Time centering is $\alpha=0$ (explicit Euler; left panel) or $\alpha=0.5$ (Crank-Nicolson; right panel). For comparison, the unit circle (blue dashed) and $\mathcal{G}(\theta,\alpha)$ with $\beta=0$ (orange solid) are also plotted.
As the right panel of Fig. 1 suggests, increasing the time-centering parameter makes the time-evolution of the error more stable; for the parameters considered the stability profile of the drift-diffusion equation is comparable to that of the static diffusion equation. Furthermore, $\alpha=0.5$ should furnish 2nd-order accuracy in time. In Section 4, we restrict to steady state for comparing accuracy of the diffusion equation to lab-frame Monte Carlo transport; thus we eliminate the issue of the time-stability considered in the present section.
4 Numerical results
In order to gauge accuracy of the drift-diffusion equation relative to transport, we consider several simple steady velocity profiles; we also provide the corresponding semi-relativistic drift-diffusion solutions to Eq. (49). In particular, in Section 4.1 we examine the solutions for several uniform constant $\beta$ profiles, removing the effect of adiabatic expansion and Doppler shift. In Section 4.2, We test a few values of the velocity gradient, for a velocity linearly proportional to $x$ (sometimes referred to as “homologous flow” in supernova literature). Given the local time derivative of velocity is 0 in these problems, in principle the spatial velocity gradient must be small to warrant application of Eq. (51). Finally, in Section 4.3 we examine a jump discontinuity that breaks velocity continuity at a spatial cell edge.
Our reference solution is a special-relativistic Monte Carlo transport implementation, where particle motion, or streaming, is performed in the lab frame while particle interactions are performed in the comoving frame. We employ a linear sub-cell profile for velocity, so the particle distance to collision is
$d_{p}=-\frac{\ln(\xi)}{(\sigma_{t,p}-\sigma_{a,p})}=-\frac{\ln(\xi)}{\gamma_{p}(1-\mu\beta_{p})(\sigma_{t,0}-\sigma_{a,0})}\;\;,$ (75)
where subscript $p$ denotes evaluation of the quantity at the location of the particle, $x_{p}$,
$\beta_{p}=(\beta_{i+1/2}^{-}-\beta_{i-1/2}^{+})\frac{(x_{p}-x_{i-1/2})}{\Delta x_{i}}+\beta_{i-1/2}^{+}\;\;,$ (76a)
$\gamma_{p}=\frac{1}{\sqrt{1-\beta_{p}^{2}}}\;\;.$ (76b)
We note that this distance formula does not take into account the change in velocity over the particle path traversing distance $d$. In order to obtain the comoving scalar intensity from the Monte Carlo, we tally a path-length estimator
$\phi_{0,i,{\rm MC}}=\frac{c}{\Delta x_{i}}\sum_{p}\sum_{d_{p}\in i}\gamma_{p}^{2}(1-\beta_{p}\mu_{p})^{2}E_{p}\frac{(1-e^{-\sigma_{a,p}d_{p}})}{\sigma_{a,p}}\;\;,$ (77)
where $d_{p}\in i$ implies the sum is over the set of tracks of particle $p$ inside cell $i$. We note that expanding the quadratic and distributing the sum furnishes path length estimators for lab-frame energy density, flux and pressure in the context of the standard energy-momentum tensor transformation for comoving energy density (see, for instance, [castor2004], Chapter 6).
For the following calculations, we fix the spatial domain $x\in[0,1]$ with 128 spatial cells. We also fix the comoving opacities as $\sigma_{t,0}=128$ and $\sigma_{a,0}=1/2$. All problems have vacuum boundary conditions on both sides of the spatial domain. Units are omitted as the linear transport and diffusion equations are spatially scale-free, so $\sigma_{t,0}=128$ and $\partial\beta/\partial x=0.03$ are 128 mean-free paths and a change in $\beta$ by 0.03 over some unit of length (e.g. the radius of a star). For the MC transport, we employ 512 particles per cell for the uniform volume source and 65,536 particles for the point source.
4.1 Constant $\beta$
Here we compare the drift-diffusion equation with lab frame MC transport at several constant velocities. This removes the effect of the adiabatic and Doppler shift term in the drift diffusion equation, isolating the effect of the $\gamma$-factors. Figure 2 has scaled comoving scalar flux versus position, where each MC (solid lines) result has its peak normalized to 1, and the corresponding diffusion results are scaled by the same resulting factor, per $\beta$ value. The result for $\beta=0$ merely confirms the static solutions are consistent. At $\beta=0.3$, we see the effect of beaming in the MC and the effect of advection in diffusion cause the solutions to skew to the right. Moreover, at $\beta=0.3$, all three results agree closely. However, at $\beta=0.6$ and 0.9, we see that the semi-relativistic result suffers in comparison to the MC result, consistent with O($\beta^{2}$) effects from the $\gamma$-factors becoming important. The fully relativistic drift-diffusion result agrees closely at each $\beta$, but is missing a spike near the source present in the MC result.
Figure 2: Scaled scalar flux versus spatial coordinate for several values of constant $\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
4.2 Homologous flow, $\beta=\beta_{c}+(x-1/2)\Delta\beta$
We next test a linear profile of the form $\beta=0.6+(x-1/2)\Delta\beta$, where $\Delta\beta=0.06$ or 0.6. Otherwise, we preserve the settings from Section 4.1. Figure 3 has comoving scaler flux versus position for the two values for MC (solid lines), semi-relativistic diffusion (dashed lines), and fully-relativistic diffusion (dotted lines). The fully relativistic solution again gives better agreement to the MC transport result. However, for $\Delta\beta=0.6$ we see some discrepancy towards the right boundary, where the fully-relativistic diffusion solution becomes closer to that of the semi-relativistic diffusion. It is possible the degree of anisotropy near the higher bound in the MC solution is not captured in the asymptotic boundary condition we have implemented.
Figure 3: Scaled scalar flux versus spatial coordinate for several values of $\Delta\beta=0.06,0.6$ in $\beta=0.6+(x-1/2)\Delta\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
4.3 Velocity jump, $\beta=\beta_{l}+\Theta(x-1/2)\Delta\beta$
Finally, we consider a problem with a very sharp, localized velocity gradient, where our velocity discretization approximates $\beta=0.3+\Theta(x-1/2)\Delta\beta$, where $\Theta(\cdot)$ is again the unit step function. Departing from the previous problems we consider a uniform comoving volume source. The jump in velocity at $x=1/2$ incurs a significant change in the comoving scalar intensity, as seen in Fig. 4 for all three solutions. In Fig. 4 we have MC transport (solid lines), semi-relativistic (dashed lines) and fully-relativistic (dotted lines) diffusion. We see that all three solutions are comparable for the $\Delta\beta=0.3$ jump (to $\beta=0.6$ at $x\geq 1/2$). For the larger jump $\Delta\beta=0.6$, we see that the fully relativistic solution is in better agreement in the higher-velocity region, but all three solutions agree fairly well in the lower-velocity region.
Figure 4: Scaled scalar flux versus spatial coordinate for several values of $\Delta\beta=0.06,0.6$ in $\beta=0.3+\Theta(x-1/2)\Delta\beta$. The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
5 Conclusions
We have derived a fully-relativistic grey diffusion equation for continuous-direction 1D planar linear radiative transfer using a lab-frame-based asymptotic analysis. This analysis makes use of a family of parameter scaling that is intended to be in proximity to the standard static-material background scaling used in linear transport theory. The resulting fully-relativistic diffusion equation has a diffusion operator that scales as $1/\gamma^{3}$, consistent with the two-direction Poisson-Kac-based derivation of [giona2017].
We have scrutinized conditions for which the asymptotic analysis gives a closed form of the diffusion equation, which are not merely the standard parameter scaling relationships in the comoving frame; these give several pathologies. One pathology is an occurance of 0 being equal to a non-zero term in the O($\varepsilon$) equation, which can be avoided by rescaling the comoving time derivative plus an anisotropic coefficient of the angular intensity. After this adjustment, another pathology is in the closed form of the equation: without asymptotically scaling the Lagrangian derivative of velocity, we see a parity violation of the solution when inverting the spatial coordinate and negating the velocity (the solution does not behave forward and backward in $x$ in the same way). Thus we must asymptotically scale the Lagrangian derivative of the velocity (or $\beta$), which indicates the resulting fully-relativistic diffusion equation can only be reliably accurate when fluid parcels are nearly not accelerating (this condition nearly exactly holds for many supernova ejecta, which follows $v=x/t$). The fully-relativistic 1D diffusion equation is straightforwardly amenable to a DDMC treatment, and hence hybridization with MC transport for acceleration.
We have also presented several simple numerical tests of the fully-relativistic diffusion equation, comparing to the $\gamma$-less semi-relativistic diffusion equation commonly used in the literature, and to a lab-frame Monte Carlo transport solution. For the problems tested, we see that the $\gamma$-dressing of the fully-relativistic diffusion equation furnishes an improvement in agreement to the lab-frame transport at $\beta\gtrsim 0.6$, relative to the semi-relativistic diffusion result. All solutions have been presented for steady-state in the lab frame, but the extension to time-dependence is amenable to standard time-discretization; however, for explicit or semi-implicit schemes care must be taken to select a sufficiently small time step size, due to the CFL-type condition imposed by the advection term in the Lagrangian time derivative.
The main objectives for future work on this topic are as follows:
  1. 1.
    thermal radiative transfer,
  2. 2.
    hybridizing with Monte Carlo transport,
  3. 3.
    extending to 3D and non-planar geometries,
  4. 4.
    and incorporating frequency dependence.
The standard static-material scaling rules for thermal radiative transfer are somewhat different than linear transport: the absorption opacity is scaled to be asymptotically large and one considers a matter equation with an asymptotically small heat capacity [larsen1983]. The time derivative is scaled as in the linear transport parameter scaling. This approach is evidently distinct from the opacity-only scaling method of [thomas1930, anderson1972] considered in the general-relativistic framework. General relativistic codes often use a time increment that does not correspond to a simple lab-frame picture, however, and the time variable is on more equal footing with the spatial variables (see Appendix Section C).
For 3D, we must consider the finding of [giona2017] that shows an anisotropic effective diffusion coefficient. While this result is still for discrete directions, we anticipate a similar effect for continuous direction in 3D. One possible path forward on a 3D derivation is to consider the analysis on a piecewise constant velocity field, where one can rotate the spatial coordinate system to have an axis aligned with the velocity in each cell; then it may be possible to follow much of the analysis given here, but with an additional step of inverting the spatial rotation (and projecting it onto a stencil that is compatible with the cells).
6 Acknowledgments
We thank Nick Gentile for useful discussions. This work has been assigned document release number LA-UR-25-31726. This work was supported by the U.S. Department of Energy through the Los Alamos National Laboratory. Los Alamos National Laboratory is operated by Triad National Security, LLC, for the National Nuclear Security Administration of U.S. Department of Energy (Contract No. 89233218CNA000001).
Appendix A $\Lambda$-integral recursion
Integrals of $\lambda_{n,k}$ over $\mu$ take the form,
$\Lambda_{n,k}=\int_{-1}^{1}\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{n}\mu^{k}d\mu=\frac{1}{\gamma^{n}}\int_{-1}^{1}\left(\frac{1}{1-\beta\mu}\right)^{n}\mu^{k}d\mu\;\;,$ (78)
where $n$ and $k$ are assumed to be non-negative integers. Substituting $y=1-\beta\mu$,
$\Lambda_{n,k}=\frac{1}{\gamma^{n}}\int_{1+\beta}^{1-\beta}\frac{1}{y^{n}}\left(\frac{1-y}{\beta}\right)^{k}\left(-\frac{1}{\beta}\right)dy=\frac{1}{\gamma^{n}\beta^{k+1}}\int_{1-\beta}^{1+\beta}\frac{1}{y^{n}}(1-y)^{k}dy\\ =\frac{1}{\gamma^{n}\beta^{k+1}}\int_{1-\beta}^{1+\beta}\frac{1}{y^{n}}\sum_{j=0}^{k}{k\choose j}(-y)^{j}dy=\frac{1}{\gamma^{n}\beta^{k+1}}\sum_{j=0}^{k}{k\choose j}(-1)^{j}\int_{1-\beta}^{1+\beta}y^{j-n}dy\;\;.$ (79)
The rightmost integral evaluates to
$\int_{1-\beta}^{1+\beta}y^{j-n}dy=\begin{cases}\displaystyle\frac{1}{j-n+1}\left((1+\beta)^{j-n+1}-(1-\beta)^{j-n+1}\right)\;\;,\;\;j\not=n-1\;\;,\\ \\ \displaystyle\ln\left(\frac{1+\beta}{1-\beta}\right)\;\;,\;\;j=n-1\;\;.\end{cases}$ (80)
If j<n1j<n-1,
$\int_{1-\beta}^{1+\beta}y^{j-n}dy=\frac{1}{n-1-j}\left(\frac{1}{(1-\beta)^{n-1-j}}-\frac{1}{(1+\beta)^{n-1-j}}\right)\\ =\frac{\gamma^{2(n-1-j)}}{n-1-j}\left((1+\beta)^{n-1-j}-(1-\beta)^{n-1-j}\right)=\frac{2\gamma^{2(n-1-j)}}{n-1-j}\sum_{l=0}^{2l+1\leq n-1-j}{n-1-j\choose 2l+1}\beta^{2l+1}\;\;.$ (81)
A corollary of Eq. (79) is $\Lambda_{n,k}$ is determined by $\Lambda_{n-j,0}$ for $j\in\{0,\ldots,k\}$; given
$\int_{1-\beta}^{1+\beta}y^{j-n}dy=\sum_{j^{\prime}=0}^{0}{0\choose j^{\prime}}(-1)^{j^{\prime}}\int_{1-\beta}^{1+\beta}y^{j^{\prime}-(n-j)}dy=\gamma^{n-j}\beta\Lambda_{n-j,0}\;\;,$ (82)
then
$\Lambda_{n,k}=\frac{1}{\gamma^{n}\beta^{k}}\sum_{j=0}^{k}{k\choose j}(-1)^{j}\gamma^{n-j}\Lambda_{n-j,0}\;\;.$ (83)
For $n$ and $k$ greater than 1,
$\Lambda_{n,k}=\frac{1}{\gamma^{n}\beta^{k}}\left(-\gamma^{n-1}\beta^{k-1}\Lambda_{n-1,k-1}+\gamma^{n}\beta^{k-1}\Lambda_{n,k-1}\right)=\frac{1}{\beta}\Lambda_{n,k-1}-\frac{1}{\gamma\beta}\Lambda_{n-1,k-1}\;\;,$ (84)
which follows from
$\int\frac{y^{k}}{(1-y)^{n}}dy=\int\frac{y^{k}-y^{k-1}}{(1-y)^{n}}dy+\int\frac{y^{k-1}}{(1-y)^{n}}dy=-\int\frac{y^{k-1}}{(1-y)^{n-1}}dy+\int\frac{y^{k-1}}{(1-y)^{n}}dy$ (85)
(it can also be derived from Eq. (83) using recursion of binomial coefficients). Figure 5 shows an example diagram of the recursion given by Eq. (84), relevant to calculations of lab-frame quantities from different $\mu_{0}$-expansions of comoving intensity, $\psi_{0}$, evaluated in the following sections.
$\Lambda_{2,0}$$\Lambda_{3,0}$$\Lambda_{4,0}$$\Lambda_{5,0}$$\Lambda_{3,1}$$\Lambda_{4,1}$$\Lambda_{5,1}$$\Lambda_{4,2}$$\Lambda_{5,2}$$\Lambda_{5,3}\leftarrow$ isotropic $\psi_{0}$anisotropic $\psi_{0}\rightarrow$
Figure 5: Example diagram of triangular recursion given by Eq. (84). The first layer of orange rectangles is evaluated with Eqs. (79) and (81). This particular diagram generates the factors needed for evaluating lab-frame $\mu$-weighted integrals from the corresponding comoving $\mu_{0}$-weighted integrals, up to linear anisotropy in the comoving frame. The red dashed line separates the factors needed for isotropic comoving intensity (left) from those needed for linear anisotropic comoving intensity (right).
Evaluating the $k=0$ integrals with Eq. (81), and using the recursion pattern in Fig. 5 gives Table 1. Notably, odd values of $k$ correspond to terms with lowest power in $\beta$ greater than 0, so that they vanish when $\beta=0$, as expected by inspection of Eq. (78).
k \ n 2 3 4 5
0 22 $2\gamma$ $\frac{2\gamma^{2}}{3}(3+\beta^{2})$ $2\gamma^{3}(1+\beta^{2})$
1 - $2\gamma\beta$ $\frac{8\gamma^{2}}{3}\beta$ $\frac{2\gamma^{3}}{3}\beta(5+\beta^{2})$
2 - - $\frac{2\gamma^{2}}{3}(1+3\beta^{2})$ $\frac{2\gamma^{3}}{3}(1+5\beta^{2})$
3 - - - $2\gamma^{3}\beta(1+\beta^{2})$
Table 1: Evaluation of recursion diagram in Fig. 5, using Eq. (81) for $k=0$ and Eq. (84) for $k>0$.
Appendix B Lab-frame P1 expansion
For convenience, we rewrite the lab-frame transport equation here:
$\frac{1}{c}\frac{\partial\psi}{\partial t}+\mu\frac{\partial\psi}{\partial x}+\sigma_{t,0}\gamma(1-\beta\mu)\psi=\frac{1}{2}\left(\frac{\sigma_{s,0}}{\gamma(1-\beta\mu)^{3}}\right)\int_{-1}^{1}(1-\beta\mu^{\prime})^{2}\psi(\mu^{\prime})d\mu^{\prime}+\frac{q}{2}\;\;.$ (86)
Defining
$E=\frac{1}{c}\int_{-1}^{1}\psi d\mu\;\;,$ (87a)
$F=\int_{-1}^{1}\mu\psi d\mu\;\;,$ (87b)
$P=\frac{1}{c}\int_{-1}^{1}\mu^{2}\psi d\mu\;\;,$ (87c)
$Q_{E}=\frac{1}{2}\int_{-1}^{1}qd\mu\;\;,$ (87d)
$Q_{M}=\frac{1}{2}\int_{-1}^{1}\mu qd\mu\;\;,$ (87e)
the 0th and 1st moment angular integrals of Eq. (86) are
$\frac{\partial E}{\partial t}+\frac{\partial F}{\partial x}+\sigma_{t,0}\gamma(cE-\beta F)=\sigma_{s,0}\gamma^{3}(cE-2\beta F+c\beta^{2}P)+Q_{E}\;\;,$ (88a)
$\frac{1}{c}\frac{\partial F}{\partial t}+c\frac{\partial P}{\partial x}+\sigma_{t,0}\gamma(F-c\beta P)=\sigma_{s,0}\gamma^{3}\beta(cE-2\beta F+c\beta^{2}P)+Q_{M}\;\;.$ (88b)
B.1 Lab frame for isotropic comoving intensity
If the comoving frequency-integrated intensity is isotropic,
$\psi_{0}=\int_{0}^{\infty}\psi_{0,\nu_{0}}d\nu_{0}=\frac{cE_{0}}{2}\;\;,$ (89)
then the lab-frame intensity satisfies
$\psi=\int_{0}^{\infty}\psi_{\nu}d\nu=\int_{0}^{\infty}\left(\frac{\nu}{\nu_{0}}\right)^{3}\psi_{0,\nu_{0}}d\nu=\int_{0}^{\infty}\left(\frac{\nu}{\nu_{0}}\right)^{3}\psi_{0,\nu_{0}}\left(\frac{\nu}{\nu_{0}}\right)d\nu_{0}\\ =\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{4}\int_{0}^{\infty}\psi_{0,\nu_{0}}d\nu_{0}=\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{4}\psi_{0}=\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{4}\frac{cE_{0}}{2}\;\;.$ (90)
The lab-frame energy density, flux, and pressure are the 0th, 1st, and 2nd moments in lab-frame angle $\mu$,
$E=\frac{1}{c}\int_{-1}^{1}\psi d\mu=\frac{E_{0}}{2}\Lambda_{4,0}\;\;,$ (91a)
$F=\int_{-1}^{1}\mu\psi d\mu=\frac{cE_{0}}{2}\Lambda_{4,1}\;\;,$ (91b)
$P=\frac{1}{c}\int_{-1}^{1}\mu^{2}\psi d\mu=\frac{E_{0}}{2}\Lambda_{4,2}\;\;.$ (91c)
Using the values from Table 1,
$E=\frac{E_{0}}{2}\Lambda_{4,0}=\gamma^{2}\left(\frac{3+\beta^{2}}{3}\right)E_{0}\;\;,$ (92a)
$F=\frac{cE_{0}}{2}\Lambda_{4,1}=\gamma^{2}\frac{4}{3}c\beta E_{0}\;\;,$ (92b)
$P=\frac{E_{0}}{2}\Lambda_{4,2}=\gamma^{2}\left(\frac{1+3\beta^{2}}{3}\right)E_{0}\;\;.$ (92c)
We can now incorporate these identities into Eq. (88)a, after simplification
$\frac{\partial}{\partial t}\left(\gamma^{2}\left(\frac{3+\beta^{2}}{3}\right)E_{0}\right)+\frac{\partial}{\partial x}\left(\gamma^{2}\frac{4}{3}c\beta E_{0}\right)+\sigma_{t,0}\gamma cE_{0}=\sigma_{s,0}\gamma cE_{0}+Q_{E}\;\;,$ (93)
which can be re-written terms of $E$, $F$, or $P$ as the only dependent variable. This equation describes the effects of advection, expansion, absorption (subtracting the scattering term from the total attenuation), and Doppler shift (in grey-integrated form), given some source $Q_{E}$. Writing in terms of $E$ and cancelling the scattering term,
$\frac{\partial E}{\partial t}+\frac{\partial}{\partial x}\left(\frac{4c\beta}{3+\beta^{2}}E\right)+\frac{1}{\gamma}\sigma_{a,0}\left(\frac{3}{3+\beta^{2}}\right)cE=Q_{E}\;\;.$ (94)
Taking the limit as $\beta\rightarrow 1$ of this equation,
$\frac{\partial E}{\partial t}+c\frac{\partial E}{\partial x}=Q_{E}\;\;.$ (95)
Thus at near-light speed, an isotropic comoving solution corresponds to a simple light-speed advection lab-frame solution, along $x$. It is also straightforward to show the lab-frame signal speed implied by Eq. (94) is causally bounded:
$\frac{4|\beta|}{3+\beta^{2}}\leq 1\iff 0\leq\beta^{2}-4|\beta|+3=(3-|\beta|)(1-|\beta|)\;\;.$
B.2 Lab frame for linearly anisotropic comoving intensity
Here we generalize Eq. (89),
$\psi_{0}=\int_{0}^{\infty}\psi_{0,\nu_{0}}d\nu_{0}=\frac{cE_{0}}{2}+\frac{3}{2}\mu_{0}F_{0}\;\;.$ (96)
Thus $F_{0}$ is the comoving flux. In 1D planar geometry the Lorentz transform of direction reduces to
$\mu_{0}=\frac{\mu-\beta}{1-\beta\mu}\;\;,$ (97)
so the lab-frame intensity is
$\psi=\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{4}\psi_{0}=\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{4}\left(\frac{cE_{0}}{2}+\frac{3}{2}\left(\frac{\mu-\beta}{1-\beta\mu}\right)F_{0}\right)\\ =\tilde{\psi}+\frac{3\gamma}{2}\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{5}\left(\mu-\beta\right)F_{0}\;\;,$ (98)
where $\tilde{\phi}$ has been introduced to account for the contribution of the comoving isotropic intensity term to the lab-frame intensity. The lab-frame energy density, flux, and pressure
$E=\tilde{E}+\frac{3\gamma}{2c}F_{0}\int_{-1}^{1}\left(\frac{1}{\gamma(1-\beta\mu)}\right)^{5}\left(\mu-\beta\right)d\mu=\tilde{E}+\frac{3\gamma}{2c}F_{0}\left(\Lambda_{5,1}-\beta\Lambda_{5,0}\right)\;\;,$ (99a)
$F=\tilde{F}+\frac{3\gamma}{2}F_{0}\left(\Lambda_{5,2}-\beta\Lambda_{5,1}\right)\;\;,$ (99b)
$P=\tilde{P}+\frac{3\gamma}{2c}F_{0}\left(\Lambda_{5,3}-\beta\Lambda_{5,2}\right)\;\;.$ (99c)
where $\tilde{E}$, $\tilde{F}$ and $\tilde{P}$ are the moments of $\tilde{\psi}$, and are given by Eq. (92). Using the rightmost column of Table 1, the $\Lambda$-coefficients in Eqs. (99) become
$\Lambda_{5,1}-\beta\Lambda_{5,0}=2\gamma^{3}\left(\frac{1}{3}\beta(5+\beta^{2})-\beta(1+\beta^{2})\right)=\frac{4\gamma^{3}}{3}\beta\left(1-\beta^{2}\right)=\frac{4\gamma}{3}\beta\;\;,$ (100a)
$\Lambda_{5,2}-\beta\Lambda_{5,1}=2\gamma^{3}\left(\frac{1}{3}\left(1+5\beta^{2}\right)-\frac{1}{3}\beta^{2}(5+\beta^{2})\right)=\frac{2\gamma^{3}}{3}\left(1+5\beta^{2}-\beta^{2}(5+\beta^{2})\right)=\frac{2\gamma^{3}}{3}\left(1-\beta^{4}\right)$
$=\frac{2\gamma}{3}(1+\beta^{2})\;\;,$ (100b)
$\Lambda_{5,3}-\beta\Lambda_{5,2}=2\gamma^{3}\left(\beta(1+\beta^{2})-\beta\frac{1}{3}\left(1+5\beta^{2}\right)\right)=\frac{2\gamma^{3}}{3}\beta\left(3+3\beta^{2}-1-5\beta^{2}\right)=\frac{4\gamma^{3}}{3}\beta\left(1-\beta^{2}\right)$
$=\frac{4\gamma}{3}\beta\;\;.$ (100c)
Evaluating Eqs. (99) with Eqs. (100)
$E=\tilde{E}+\frac{3\gamma}{2c}F_{0}\left(\Lambda_{5,1}-\beta\Lambda_{5,0}\right)=\tilde{E}+2\gamma^{2}\frac{1}{c}F_{0}\beta=\gamma^{2}\left(\left(\frac{3+\beta^{2}}{3}\right)E_{0}+2\frac{\beta}{c}F_{0}\right)\;\;,$ (101a)
$F=\tilde{F}+\frac{3\gamma}{2}F_{0}\left(\Lambda_{5,2}-\beta\Lambda_{5,1}\right)=\tilde{F}+\gamma^{2}F_{0}(1+\beta^{2})=\gamma^{2}\left(\frac{4}{3}c\beta E_{0}+F_{0}(1+\beta^{2})\right)\;\;,$ (101b)
$P=\tilde{P}+\frac{3\gamma}{2c}F_{0}\left(\Lambda_{5,3}-\beta\Lambda_{5,2}\right)=\tilde{P}+2\gamma^{2}\frac{1}{c}F_{0}\beta=\gamma^{2}\left(\left(\frac{1+3\beta^{2}}{3}\right)E_{0}+2\frac{\beta}{c}F_{0}\right)\;\;.$ (101c)
We see that for $\beta=0$, $F=F_{0}$, as expected. The linear aniostropic contribution to comoving intensity adds the same contribution to $\tilde{E}$ and $\tilde{P}$ to obtain $E$ and $P$, respectively. We may write the equations for $E$ and $F$ as
$\left(\begin{array}[]{c}E\\ F\end{array}\right)=\gamma^{2}\left(\begin{array}[]{cc}\frac{3+\beta^{2}}{3}&2\frac{\beta}{c}\\ \frac{4}{3}c\beta&1+\beta^{2}\end{array}\right)\left(\begin{array}[]{c}E_{0}\\ F_{0}\end{array}\right)\;\;.$ (102)
Inverting the 2x2 matrix,
$\left(\begin{array}[]{c}E_{0}\\ F_{0}\end{array}\right)=\frac{3}{\gamma^{2}(3-4\beta^{2}+\beta^{4})}\left(\begin{array}[]{cc}1+\beta^{2}&-2\frac{\beta}{c}\\ -\frac{4}{3}c\beta&\frac{3+\beta^{2}}{3}\end{array}\right)\left(\begin{array}[]{c}E\\ F\end{array}\right)=\frac{3}{(3-\beta^{2})}\left(\begin{array}[]{cc}1+\beta^{2}&-2\frac{\beta}{c}\\ -\frac{4}{3}c\beta&\frac{3+\beta^{2}}{3}\end{array}\right)\left(\begin{array}[]{c}E\\ F\end{array}\right)$ (103)
It can be seen that $\beta=0$ reduces the 2x2 matrix to the identity matrix, in Eqs. (102) and (103). Using Eq. (103), $P$ can be expressed in terms of $E$ and $F$,
$P=\gamma^{2}\frac{3}{(3-\beta^{2})}\left(\left(\frac{1+3\beta^{2}}{3}\right)\left((1+\beta^{2})E-2\frac{\beta}{c}F\right)+2\frac{\beta}{c}\left(\frac{3+\beta^{2}}{3}F-\frac{4}{3}c\beta E\right)\right)\\ =\frac{\gamma^{2}}{(3-\beta^{2})}\left(((1+3\beta^{2})(1+\beta^{2})-8\beta^{2})E+((3+\beta^{2})-(1+3\beta^{2}))2\frac{\beta}{c}F\right)\\ =\frac{\gamma^{2}}{(3-\beta^{2})}\left((1-\beta^{2})(1-3\beta^{2})E+4(1-\beta^{2})\frac{\beta}{c}F\right)=\frac{1}{(3-\beta^{2})}\left((1-3\beta^{2})E+4\frac{\beta}{c}F\right)$ (104)
Equation (104) has the following limits in $\beta$,
$\lim_{\beta\rightarrow 0}P=\frac{E}{3}\;\;,$ (105a)
$\lim_{\beta\rightarrow 1}P=\frac{2}{c}F-E\;\;,$ (105b)
where the $\beta\rightarrow 0$ limit is the expected static isotropic pressure. The $\beta\rightarrow 1$ limit is consistent with (but not derived from) $\psi=cE\delta(\mu-1)$, where $\delta(\cdot)$ is the Dirac delta distribution.
Augmenting Eqs. (88) with Eq. (104), the closed system of equations is
$\frac{\partial E}{\partial t}+\frac{\partial F}{\partial x}+\sigma_{t,0}\gamma(cE-\beta F)=\sigma_{s,0}\gamma^{3}(cE-2\beta F+c\beta^{2}P)+Q_{E}\;\;,$ (106a)
$\frac{1}{c}\frac{\partial F}{\partial t}+c\frac{\partial P}{\partial x}+\sigma_{t,0}\gamma(F-c\beta P)=\sigma_{s,0}\gamma^{3}\beta(cE-2\beta F+c\beta^{2}P)+Q_{M}\;\;,$ (106b)
$P=\frac{1}{(3-\beta^{2})}\left((1-3\beta^{2})E+4\frac{\beta}{c}F\right)\;\;.$ (106c)
B.3 O($\beta$) and neglecting derivatives of $F$ and $\beta F$
B.3.1 O($\beta$)
To O($\beta$), Eqs. (106) become
$\frac{\partial E}{\partial t}+\frac{\partial F}{\partial x}+\sigma_{t,0}(cE-\beta F)=\sigma_{s,0}(cE-2\beta F)+Q_{E}\;\;,$ (107a)
$\frac{1}{c}\frac{\partial F}{\partial t}+c\frac{\partial P}{\partial x}+\sigma_{t,0}(F-c\beta P)=c\sigma_{s,0}\beta E+Q_{M}\;\;,$ (107b)
$P=\frac{1}{3}\left(E+4\frac{\beta}{c}F\right)\;\;.$ (107c)
Incorporating Eq. (107)c into Eq. (107)b and neglecting O($\beta^{2}$),
$\frac{1}{c}\frac{\partial F}{\partial t}+\frac{4}{3}\frac{\partial}{\partial x}(\beta F)+\frac{c}{3}\frac{\partial E}{\partial x}+\sigma_{t,0}F-\frac{1}{3}c\beta\sigma_{t,0}E=c\beta\sigma_{s,0}E+Q_{M}\;\;.$ (108)
This can be re-written in terms of the comoving time derivative (at O($\beta$) now equivalent to the Lagrangian time derivative),
$\frac{1}{c}\frac{\partial F}{\partial t_{0}}+F\frac{\partial\beta}{\partial x}+\frac{1}{3}\frac{\partial}{\partial x}(\beta F)+\frac{c}{3}\frac{\partial E}{\partial x}+\sigma_{t,0}F-\frac{1}{3}c\beta\sigma_{t,0}E=c\beta\sigma_{s,0}E+Q_{M}\;\;.$ (109)
B.3.2 Neglecting the comoving time derivative of $F$
If we neglect the comoving time derivative of $F$ and simplify,
$\left(\frac{4}{3}\frac{\partial\beta}{\partial x}+\sigma_{t,0}\right)F+\frac{\beta}{3}\frac{\partial F}{\partial x}=-\frac{c}{3}\frac{\partial E}{\partial x}+c\beta\left(\frac{1}{3}\sigma_{t,0}+\sigma_{s,0}\right)E+Q_{M}\;\;.$ (110)
Equation (110) is a 1st order ordinary differential equation for $F$. Assuming $\beta\not=0$, the solution is
$F(x)=\left(\frac{\beta(x)}{\beta(0)}\right)^{4}e^{-3\int_{0}^{x}\sigma_{t,0}/\beta(x^{\prime})dx^{\prime}}F(0)\\ +\int_{0}^{x}\left(\frac{\beta(x^{\prime})}{\beta(0)}\right)^{4}e^{-3\int_{x^{\prime}}^{x}\sigma_{t,0}/\beta(x^{\prime\prime})dx^{\prime\prime}}\left(-\frac{c}{\beta(x^{\prime})}\frac{\partial E}{\partial x^{\prime}}+c(\sigma_{t,0}+3\sigma_{s,0})E(x^{\prime})+\frac{3}{\beta(x^{\prime})}Q_{M}(x^{\prime})\right)dx^{\prime}$ (111)
where we have taken $x=0$ to be the integration bound. If we take $\beta$ to be constant, then take the limit as $\beta$ goes to 0, then in the second term on the right side of Eq. (111) we have indeterminate forms
$\lim_{\beta\rightarrow 0}\frac{e^{-3\sigma_{t,0}(x-x^{\prime})/\beta}}{\beta}=\frac{1}{3\sigma_{t,0}}\lim_{\beta\rightarrow 0}\frac{3\sigma_{t,0}e^{-3\sigma_{t,0}(x-x^{\prime})/\beta}}{\beta}=\frac{1}{3\sigma_{t,0}}\delta^{(+)}(x-x^{\prime})\;\;,$ (112)
multiplying the gradient of energy density and $Q_{M}$, where we have introduced a half-space Dirac delta distribution in the final equality, given the indeterminate form satisfies the criteria of being a nascent half-space Dirac delta distribution (compact support and unit integral). Thus in this limit Eq. (111) reduces to
$F=-\frac{c}{3\sigma_{t,0}}\frac{\partial E}{\partial x}+\frac{1}{\sigma_{t,0}}Q_{M}\;\;,$ (113)
which is the static form of Fick’s Law (if the transport source term is isotropic in the lab frame, then $Q_{M}=0$).
B.3.3 Neglecting terms $\sim\partial(\beta F)/\partial x$
If the terms consisting of a spatial gradient of $F$ or $\beta$ multiplied by $\beta$ or $F$ are neglected, Eq. (110) reduces to
$F=-\frac{c}{3\sigma_{t,0}}\frac{\partial E}{\partial x}+c\beta\left(\frac{1}{3}+\frac{\sigma_{s,0}}{\sigma_{t,0}}\right)E+\frac{1}{\sigma_{t,0}}Q_{M}\;\;.$ (114)
Incorporating Eq. (114) into Eq. (107) and keeping O($\beta$) terms
$\frac{\partial E}{\partial t}+c\beta\frac{\partial E}{\partial x}-\frac{\partial}{\partial x}\left(\frac{c}{3\sigma_{t,0}}\frac{\partial E}{\partial x}\right)+c\frac{\partial}{\partial x}\left(\left(\frac{1}{3}+\frac{\sigma_{s,0}}{\sigma_{t,0}}\right)\beta E\right)-\frac{2}{3}c\left(1+\frac{\sigma_{s,0}}{\sigma_{t,0}}\right)\beta\frac{\partial E}{\partial x}+\sigma_{a,0}cE\\ =Q_{E}+\left(1-\frac{2\sigma_{s,0}}{\sigma_{t,0}}\right)\beta Q_{M}-\frac{\partial}{\partial x}\left(\frac{1}{\sigma_{t,0}}Q_{M}\right)\;\;.$ (115)
The comoving scattering ratio appears in several terms in Eq. (115). Taking the limit as $\sigma_{s,0}\rightarrow\sigma_{t,0}$ and simplifying,
$\frac{\partial E}{\partial t}+c\beta\frac{\partial E}{\partial x}-\frac{\partial}{\partial x}\left(\frac{c}{3\sigma_{t,0}}\frac{\partial E}{\partial x}\right)+\frac{4}{3}cE\frac{\partial\beta}{\partial x}+\sigma_{a,0}cE=Q_{E}-\beta Q_{M}-\frac{\partial}{\partial x}\left(\frac{1}{\sigma_{t,0}}Q_{M}\right)\;\;.$ (116)
Equation (116) is the 1D planar form of Eq. 6.51 of [castor2004], but with the following modifications:
  • 1.
    the gradient of $F$ is replaced by the diffusion operator on $E$ (Fick’s Law),
  • 2.
    the pressure tensor is replaced by the identity matrix multiplied by $E/3$,
  • 3.
    an additional term, $\partial(Q_{M}/\sigma_{t,0})/\partial x$, appears on the right side.
However, Eq. 6.51 of [castor2004] is for comoving $E$ and $F$. Similarly, this is the limit we obtain of (48) ((49)) in Section 2.4, but with lab-frame $E$.
Appendix C The Thomas scaling in non-relativistic 1D planar geometry
Here we briefly consider the Thomas scaling [thomas1930], as presented by [anderson1972], in a non-relativistic 1D context with constant pure absorption opacity. To the best of our understanding, the only small parameter is the mean-free path in these approaches. This should constitute a simplification of Section IV of [anderson1972]. Using similar notation to [anderson1972], the transport equation is then
$\frac{1}{\sigma}\frac{d\psi}{d\lambda}=\frac{S}{2}-\psi\;\;,$ (117)
where we have written the partial derivatives of the streaming operator in terms of the characteristic, or affine coordinate $\lambda$. Following [anderson1972], we write the asymptotic expansion directly in powers of $\sigma$.
$\psi=\sum_{k=0}^{\infty}\frac{\tilde{\psi}^{(k)}}{\sigma^{k}}\;\;.$ (118)
We note that this is equivalent to Eq. (25), by introducing a length scale $L$, and setting $\varepsilon=1/L\sigma$ and $\psi^{(k)}=L^{k}\tilde{\psi}^{(k)}$. Again following [anderson1972], matching orders to O($1/\sigma$) gives
$\tilde{\psi}^{(0)}=\frac{S}{2}\;\;,$ (119a)
$\tilde{\psi}^{(1)}=-\frac{d\psi^{(0)}}{ds}=-\frac{1}{2}\frac{dS}{d\lambda}$ (119b)
(corresponding to their equation 37 in Section IV). Applying the expansion to O($1/\sigma$) in Eq. (117) and simplifying gives
$\frac{1}{\sigma}\frac{d^{2}S}{d\lambda^{2}}=0\;\;.$ (120)
Expanding the affine derivative using the 1D planar assumption and integrating over $\mu$,
$\frac{1}{c\sigma}\frac{d^{2}S}{dt^{2}}+\frac{1}{3\sigma}\frac{d^{2}S}{dx^{2}}=0\;\;,$ (121)
where we have assumed $S$ is isotropic.
If instead we only expand the left side of Eq. (117), then relabel $S$ as $\psi$ on the left side, we obtain
$\frac{d}{d\lambda}\left(\psi-\frac{1}{\sigma}\frac{d\psi}{d\lambda}\right)=\sigma\left(\frac{S}{2}-\psi\right)\;\;,$ (122)
which upon expansion into $(x,t)$ and integration over $\mu$ gives
$\frac{1}{c}\frac{\partial\phi}{\partial t}-\frac{1}{\sigma}\left(\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}+\frac{1}{3}\frac{\partial^{2}}{\partial x^{2}}\right)\phi=\sigma(S-\phi)\;\;,$ (123)
which is nearly the Telegrapher’s equation (it is so with an imaginary relaxation coefficient of the 2nd derivative in time). Putting into the 4-vector notation of [anderson1972, shibata2011], Eq. (123) becomes
$u^{\alpha}\nabla_{\alpha}\phi-\frac{1}{3\sigma}\nabla_{\alpha}\left(\left(3u^{\alpha}u^{\delta}+h^{\alpha\delta}\right)\nabla_{\gamma}\phi\right)=\sigma(S-\phi)\;\;,$ (124)
where $\alpha$ and $\delta$ here are 4-indexes, $u^{\alpha}=(1,0,0,0)$ is the 4-velocity, and $h^{\alpha\delta}=u^{\alpha}u^{\delta}-\eta^{\alpha\delta}$ is the rank-4 projection operator, as in [anderson1972, shibata2011] ($\eta^{\alpha\delta}$ here is the Minkowski metric).
In contrast to equation 5.21 of [shibata2011], Eq. (124) has an additional time-like component in the 2nd derivative, which can be seen as a contribution of $\tilde{\phi}^{(1)}$ (consistent with equation 5.13 of [shibata2011]). Excising this time-like component would furnish a non-equilibrium diffusion equation from Eqs. (123) and (124).

### v2

@@
 It is well known to be a limiting behavior of the more general phenomenon of radiative transfer.
-For instance, photon propagation through expanding supernova ejecta can be characterized with a diffusion solution (see, for instance, [arnett1982, pinto2000]).
+For instance, photon propagation through expanding supernova ejecta can be characterized with a diffusion solution (see, for instance, [4, 29]).
 Numerical implementation of the diffusion equation can be used to accelerate radiative transfer in optically thick regions of space and frequency.
-One example is Monte Carlo (MC) accelerated with random walk (RW) or discrete diffusion Monte Carlo (DDMC) [fleck1984, gentile2001, densmore2007, densmore2012, cleveland2014, keady2017, smith2020].
-These RW and DDMC methods have been developed for applications to: neutrinos in core collapse supernova engines [abdikamalov2012], cosmic rays [harding2016, fitzaxen2021], radiation pressure feedback in stellar cluster dynamics [tsang2018], supernova and kilonova spectral synthesis [wollaeger2013, wollaeger2014, wagle2023], and Lyman $\alpha$ photon transfer during galaxy evolution [smith2018].
-While diffusion theory has found practical use in describing photon and neutrino fields in astrophysical plasmas, the infinite signal speed of the diffusion equation is incompatible with special relativity (see, for instance, [dunkel2007, gzyl2025]).
+One example is Monte Carlo (MC) accelerated with random walk (RW) or discrete diffusion Monte Carlo (DDMC) [12, 13, 8, 7, 6, 21, 32].
+These RW and DDMC methods have been developed for applications to: neutrinos in core collapse supernova engines [1], cosmic rays [18, 11], radiation pressure feedback in stellar cluster dynamics [36], supernova and kilonova spectral synthesis [38, 39, 37], and Lyman $\alpha$ photon transfer during galaxy evolution [33].
+While diffusion theory has found practical use in describing photon and neutrino fields in astrophysical plasmas, the infinite signal speed of the diffusion equation is incompatible with special relativity (see, for instance, [10, 15]).
 This can readily be seen for static material backgrounds, where the Green’s function conforms to a Gaussian distribution, which has non-zero values everywhere.
-As a result, the applicability of diffusion is limited to high-optical depth regions, unless the equation is modified to restrict the signal speed (for instance $P_{1}$, $M_{1}$ [levermore1984] and flux-limit diffusion [levermore1981]).
-We may also see the effect of Lorentz transforming the space-time derivatives in the diffusion equation.
-Suppose that a frame of reference, denoted by subscript 0, possesses particles that adhere to a 1D diffusion equation,
+As a result, the applicability of diffusion is limited to high-optical depth regions, unless the equation is modified to restrict the signal speed (for instance $P_{1}$, $M_{1}$ [24] and flux-limit diffusion [25]).
+It is straightforward to show that the diffusion equation is not Lorentz-invariant. For instance, consider the 1D planar diffusion equation,
 $\frac{\partial\phi_{0}}{\partial t_{0}}-D\frac{\partial^{2}\phi_{0}}{\partial x_{0}^{2}}=0\;\;,$
 where $t_{0}$ is time, $x_{0}$ is the spatial coordinate, $D$ is a constant diffusion coefficient, and $\phi_{0}$ is the particle density.
-Applying the standard Lorentz transformations for the partial derivatives (see, for instance, [lowrie2023]) gives
+Applying the standard Lorentz transformations for the partial derivatives (see, for instance, [27]) gives
 $\gamma\left(\frac{\partial}{\partial t}+v\frac{\partial}{\partial x}\right)\phi_{0}-D\left[\gamma\left(\frac{\partial}{\partial x}+\frac{v}{c^{2}}\frac{\partial}{\partial t}\right)\right]^{2}\phi_{0}=0\;\;,$
@@
 $\left(\frac{\partial}{\partial t}+c\beta\frac{\partial}{\partial x}\right)\phi_{0}-\gamma D\left(\frac{\partial^{2}}{\partial x^{2}}+2\frac{\beta}{c}\frac{\partial^{2}}{\partial t\partial x}+\frac{\beta^{2}}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}\right)\phi_{0}=0\;\;.$
-Thus in the new frame, distinguished here by the absence of a subscript, a second derivative in time is manifest.
-Similar to the infinitely extended signal in space, this second derivative in time must couple all time from the past to the future and vice versa.
-Consequently, this equation is not immediately amenable to numerical schemes that can only advance forward in time in the frame of $(x,t)$.
+This resembles a standard drift-diffusion equation if the second derivatives involving $t$ are neglected.
 The problem of unifying diffusion with the principles of special relativity has been recognized in some form for approximately one century; some approaches to combining them include: (i) statistical analysis of stochastic kinematics and (ii) asymptotic analysis of the radiative transfer equation with relativistic corrections.
 The latter approach is often pursued in transport theory and astrophysics literature, where the effects of collective fluid motion on particle distributions is a focus.
-Stochastic analyses originated by [hakim1965] and [dudley1966] (and further developed in subsequent works) demonstrate that relativity breaks Markovian processes in space, where one would expect that the set of spatial positions for the current particle state completely determine the next state [dudley1966, dunkel2007, giona2017].
-In developing the relativistic theory, one might first consider bounding the particle speed in a Poisson-Kac process [kac1974], which corresponds to the Telegrapher ($P_{1}$) equation and is non-Markovian [kac1974, dunkel2007].
-Spatially non-Markovian stochastic processes have been formulated in special [dunkel2007, serva2020] and general [herrmann2010] relativistic settings.
+Stochastic analyses originated by [17] and [9] (and further developed in subsequent works) demonstrate that relativity breaks Markovian processes in space, where one would expect that the set of spatial positions for the current particle state completely determine the next state [9, 10, 14].
+In developing the relativistic theory, one might first consider bounding the particle speed in a Poisson-Kac process [20], which corresponds to the Telegrapher ($P_{1}$) equation and is non-Markovian [20, 10].
+Spatially non-Markovian stochastic processes have been formulated in special [10, 30] and general [19] relativistic settings.
 Properly accounting for relativity in a collective sense, where a frame is moving relative to the frame in which diffusion is a good description, is a somewhat distinct problem.
-To this end, [giona2017] examine the particle motion as a Poisson-Kac process along discrete directions, where the process exists in its usual form in one inertial (comoving) frame, finding in 1D that the effective diffusion coefficient in another inertial frame is $1/\gamma^{3}$ times the comoving coefficient; [giona2017] observe that this follows the intuition that $D\sim\delta x^{2}/\delta t\sim(\delta x_{0}/\gamma)^{2}/(\gamma\delta t_{0})$, from Lorentz contraction of space and time dilation.
-For asymptotic analysis, the principal assumption is typically that the inverse of the macroscopic opacity is large compared to all length scales of interest [thomas1930].
+To this end, [14] examine the particle motion as a Poisson-Kac process along discrete directions, where the process exists in its usual form in one inertial (comoving) frame, finding in 1D that the effective diffusion coefficient in another inertial frame is $1/\gamma^{3}$ times the comoving coefficient; [14] observe that this follows the intuition that $D\sim\delta x^{2}/\delta t\sim(\delta x_{0}/\gamma)^{2}/(\gamma\delta t_{0})$, from Lorentz contraction of space and time dilation.
+For asymptotic analysis, the principal assumption is typically that the inverse of the macroscopic opacity is large compared to all length scales of interest [34].
 Unlike the works above that focus on the statistical mechanics of the particles, this approach usually: starts from the full radiative transfer equation, imposes the assumed parameter scalings, and expands the radiation intensity in powers of a small-scale parameter (sometimes labeled $\varepsilon$).
-Asymptotic analysis has been applied to deriving a semi-relativistic diffusion equation [morel2006, lowrie2014], where $\beta$ is assumed to be asymptotically small (equivalently: the hydrodynamic flow time scale is assumed much longer than the light crossing time at the length scale of interest).
-In the fully relativistic context, including general relativity, asymptotic analysis has been examined early on by [thomas1930] and subsequently generalized by [anderson1972] (see also [achterberg2018]).
-Building off of these works, [shibata2011] demonstrate the truncated Thorne moment formalism (see [thorne1981]) converges to a diffusion equation with modifications from fluid acceleration and shear in the optically thick limit.
-The analyses of [anderson1972] and [shibata2011] consider both asymptotically large absorption and scattering opacity.
-This is consistent with the static-material scaling that furnishes equilibrium diffusion from non-linear thermal radiative transfer, but departs in not assuming that the light crossing distance on the time scale of interest is large compared to the spatial domain of interest [larsen1983].
+Asymptotic analysis has been applied to deriving a semi-relativistic diffusion equation [28, 26], where $\beta$ is assumed to be asymptotically small (equivalently: the hydrodynamic flow time scale is assumed much longer than the light crossing time at the length scale of interest).
+In the fully relativistic context, including general relativity, asymptotic analysis has been examined early on by [34] and subsequently generalized by [3] (see also [2]).
+Building off of these works, [31] demonstrate the truncated Thorne moment formalism (see [35]) converges to a diffusion equation with modifications from fluid acceleration and shear in the optically thick limit.
+The analyses of [3] and [31] consider both asymptotically large absorption and scattering opacity.
+This is consistent with the static-material scaling that furnishes equilibrium diffusion from non-linear thermal radiative transfer, but departs in not assuming that the light crossing distance on the time scale of interest is large compared to the spatial domain of interest [23].
 In the present study, we instead consider an asymptotic analysis that restricts to a special relativistic, 1D-planar, linear transport context, and we focus on the correspondence of the lab frame (some inertial frame) to the non-inertial frame comoving with fluid parcels of the background matter.
-In a departure from the work of [anderson1972] and [shibata2011] we attempt to use the standard parameter scaling for linear transport (see, for instance, [habetler1975, larsen1992]) in the comoving frame.
+In a departure from the work of [3] and [31] we attempt to use the standard parameter scaling for linear transport (see, for instance, [16, 22]) in the comoving frame.
 Thus the scattering opacity is assumed to be asymptotically large while the absorption opacity is assumed to be asymptotically small, and there is not an equilibrium solution (for instance, the Planck or Fermi-Dirac distributions in the comoving frame).
-In doing so, our objective is to: (i) highlight apparent non-physical pathologies that occur from applying this standard scaling when velocity gradients are present, (ii) demonstrate that a particular scaling furnishes the Poisson-Kac-derived $\gamma$-dressed result of [giona2017], but for continuous directions and including the standard grey contribution from adiabatic expansion and Doppler shift (see, for instance [castor2004]), (iii) compare solutions of the diffusion equation to a relativistic linear transport solution and better-known semi-relativistic diffusion equation [castor2004].
+In doing so, our objective is to: (i) highlight apparent non-physical pathologies that occur from applying this standard scaling when velocity gradients are present, (ii) demonstrate that a particular scaling furnishes the Poisson-Kac-derived $\gamma$-dressed result of [14], but for continuous directions and including the standard grey contribution from adiabatic expansion and Doppler shift (see, for instance [5]), (iii) compare solutions of the diffusion equation to a relativistic linear transport solution and better-known semi-relativistic diffusion equation [5].
 This paper is organized as follows.
@@
 We then compare this discretization to lab-frame MC transport to provide a preliminary exploration of where the fully relativistic diffusion equation is applicable in Section 4.
-In the appendix, we provide: (A) details of the integrated Doppler function recursion, (B) the aforementioned harmonic ($P_{1}$) expansion, and finally (C) a brief non-relativistic attempt of the apparent opacity-only scaling of [thomas1930, anderson1972, shibata2011].
+In the appendix, we provide: (A) details of the integrated Doppler function recursion, (B) the aforementioned harmonic ($P_{1}$) expansion, and finally (C) a brief non-relativistic attempt of the apparent opacity-only scaling of [34, 3, 31].
 2 Derivation of 1D planar fully relativistic grey diffusion
@@
 $A[\psi]=\frac{1}{c}\frac{\partial\psi}{\partial t}=\frac{1}{c}\frac{\partial\psi}{\partial t}+\beta\frac{\partial\psi}{\partial x}+B(\mu)\psi-\beta\frac{\partial\psi}{\partial x}-B(\mu)\psi\\ \rightarrow\varepsilon\left(\frac{1}{c}\frac{\partial\psi}{\partial t}+\beta\frac{\partial\psi}{\partial x}+B(\mu)\psi\right)-\beta\frac{\partial\psi}{\partial x}-B(\mu)\psi=A_{\varepsilon}[\psi]\;\;,$ (27)
-where $A[\cdot]$ is the original time-derivative operator (or the remainder of the transport equation), $B(\nu)$ is an as-yet unknown coefficient and $A_{\varepsilon}[\cdot]$ is a $\varepsilon$-scaled form of this operator.
+where $A[\cdot]$ is the original time-derivative operator (or the remainder of the transport equation), $B(\mu)$ is an as-yet unknown coefficient and $A_{\varepsilon}[\cdot]$ is a $\varepsilon$-scaled form of this operator.
 This scaling assumes that the residual excluded from the $\varepsilon$ coefficient scales as a leading-order contribution.
@@
 Furthermore, we note that we have not assumed a constraint on the value of $\beta$ itself.
-The diffusion stencil has three factors of $1/\gamma$, consistent with the finding of [giona2017] for bi-directional 1D planar random walk.
+The diffusion stencil has three factors of $1/\gamma$, consistent with the finding of [14] for bi-directional 1D planar random walk.
 However, we have an additional $\beta$-dependent coefficient of $\phi_{0}^{(0)}$, which accounts for the energetic effect of Doppler shift and adiabatic expansion.
@@
 $\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3}\frac{\partial}{\partial x}\left(\frac{1}{\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\sigma_{a,0}\phi_{0}^{(0)}+\frac{1}{2}\int_{-1}^{1}q_{0}(\mu_{0})d\mu_{0}\;\;,$ (49)
-which is the standard semi-relativistic form (see, for instance, [castor2004], Chapter 6). If $\beta$ is constant and approaches 1, assuming $q$ and $\phi_{0}^{(0)}$ remain bounded,
+which is the standard semi-relativistic form (see, for instance, [5], Chapter 6). If $\beta$ is constant and approaches 1, assuming $q$ and $\phi_{0}^{(0)}$ remain bounded,
 $\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}=0\;\;,$ (50)
-which is the limiting behavior observed by [giona2017].
+which is the limiting behavior observed by [14].
 2.4.3 Scaling the Lagrangian time derivative of $\beta$
@@
 Incorporating Eq. (51) into the suite of scaling conditions removes the term corresponding to the Lagrangian derivative of velocity in the O($\varepsilon^{2}$) equation, resulting in
-$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3\gamma^{2}}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\frac{\sigma_{a,0}}{\gamma}\phi_{0}^{(0)}+\frac{Q_{0}}{\gamma}\;\;.$ (52)
-where we have replaced the integral of the comoving external source with $Q_{0}$.
+$\left(\frac{1}{c}\frac{\partial}{\partial t}+\beta\frac{\partial}{\partial x}\right)\phi_{0}^{(0)}+\frac{4}{3}\frac{\partial\beta}{\partial x}\phi_{0}^{(0)}-\frac{1}{3\gamma^{2}}\frac{\partial}{\partial x}\left(\frac{1}{\gamma\sigma_{t,0}}\frac{\partial\phi_{0}^{(0)}}{\partial x}\right)=-\frac{\sigma_{a,0}}{\gamma}\phi_{0}^{(0)}+\frac{Q_{0}}{\gamma}\;\;,$ (52)
+where
+$Q_{0}\equiv\frac{1}{2}\int_{-1}^{1}q_{0}(\mu_{0})d\mu_{0}\;\;.$
 This equation now has the same coefficient for adiabatic and expansion effects as the standard semi-relativistic form, so retains mirrored behavior in scalar intensity when $(x,\beta)\rightarrow(-x,-\beta)$.
@@
 (52) we use to explore properties of the solutions in Section 4.
-We solve the resulting discretized system deterministically, but we note that there exists a straightforward Discrete Diffusion Monte Carlo (DDMC) [gentile2001, densmore2007, abdikamalov2012] interpretation of the discretized terms, and write them following the notation of [densmore2007].
+We solve the resulting discretized system deterministically, but we note that there exists a straightforward Discrete Diffusion Monte Carlo (DDMC) [13, 8, 1] interpretation of the discretized terms, and write them following the notation of [8].
 In 1D, a grey deterministic solution is straightforward to implement, and we do not seek to hybridize lab-frame MC transport with DDMC in the scope of this work.
@@
 To simplify the presentation of the spatial discretization that follows, we introduce a label for the coefficient of $\phi_{0}^{(0)}$ on the left side,
-$\eta_{i}=\frac{4}{3}\left.\frac{\partial\beta}{\partial x}\right|_{i}$ (54)
+$\eta_{i}\equiv\frac{4}{3}\left.\frac{\partial\beta}{\partial x}\right|_{i}$ (54)
 where subscript $i$ is the index of the spatial cell, indicating evaluation at the cell.
@@
 $\sigma_{A,i\rightarrow i+1}=\frac{1}{\Delta x_{i}}\Theta(\beta_{i+1/2}^{-})\beta_{i+1/2}^{-}\;\;,$ (59b)
-can be viewed as advection “leakage opacities” [densmore2007] (see [abdikamalov2012] for an operator-split continuous treatment of advection).
-The diffusion operator terms can be evaluated in the usual way in the domain interior [densmore2007],
+can be viewed as advection “leakage opacities” [8] (see [1] for an operator-split continuous treatment of advection).
+The diffusion operator terms can be evaluated in the usual way in the domain interior [8],
 $\frac{1}{3\gamma_{i+1/2}\sigma_{t,0,i+1/2}}\left.\frac{\partial\phi_{0}}{\partial x}\right|_{i+1/2}=\frac{2(\phi_{0,i+1}-\phi_{0,i})}{3(\gamma_{i+1/2}^{-}\sigma_{t,0,i+1/2}^{-}\Delta x_{i}+\gamma_{i+1/2}^{+}\sigma_{t,0,i+1/2}^{+}\Delta x_{i+1})}\;\;,$ (60)
@@
 The formulation permits several options for discretizing the velocity field, for instance: if $\gamma$ is continuous then $\gamma_{i+1/2}^{-}=\gamma_{i+1/2}^{+}=\gamma_{i+1/2}$; if $\gamma$ is piecewise-constant, then one may set $\gamma_{i+1/2}^{-}=\gamma_{i}$ and $\gamma_{i+1/2}^{+}=\gamma_{i+1}$.
-Consequently, the diffusion leakage opacities, corresponding to the standard non-relativistic form [densmore2007], are
+Consequently, the diffusion leakage opacities, corresponding to the standard non-relativistic form [8], are
 $\sigma_{D,i\rightarrow i-1}=\frac{2}{3\gamma_{i}^{2}\Delta x_{i}}\left(\frac{1}{\gamma_{i-1/2}^{+}\sigma_{t,0,i-1/2}^{+}\Delta x_{i}+\gamma_{i-1/2}^{-}\sigma_{t,0,i-1/2}^{-}\Delta x_{i-1}}\right)\;\;,$ (61a)
@@
 Equation (62) is very similar to a standard grey diffusion stencil, but with the addition of advection leakage, the $\eta$ coefficient, and $\gamma$ factors multiplying diffusion leakage opacities.
-Finally, to treat the boundary, we assume the asymptotic boundary layer condition used by [densmore2007] holds in the comoving frame,
+Finally, to treat the boundary, we assume the asymptotic boundary layer condition used by [8] holds in the comoving frame,
 $2\int_{0}^{1}W(\mu_{0})\psi_{0}^{-}(x_{i-1/2},\mu_{0})d\mu_{0}=\phi_{0}(x_{i-1/2})-\frac{\lambda}{\gamma\sigma_{t,0,i-1/2}}\left.\frac{\partial\phi_{0}}{\partial x}\right|_{i-1/2}\;\;,$ (63)
-where $W(\cdot)$ is the angular weighting function for intensity incident just below $x_{i-1/2}$, $\psi_{0}^{-}$, and $\lambda$ is the extrapolation distance [habetler1975].
+where $W(\cdot)$ is the angular weighting function for intensity incident just below $x_{i-1/2}$, $\psi_{0}^{-}$, and $\lambda$ is the extrapolation distance [16].
 Equation (63) is consistent with applying scaling Eqs.
-(47) to the transport equation, keeping terms up to O($\varepsilon$), transforming intensity and opacity to the comoving frame, and then applying the boundary layer analysis of [habetler1975] (the factor of $\gamma$ from the transformation of $\sigma_{t}$ to $\sigma_{t,0}$ can be included in the transformation of the spatial coordinate to the stretched optical depth parameter).
+(53) to the transport equation, keeping terms up to O($\varepsilon$), transforming intensity and opacity to the comoving frame, and then applying the boundary layer analysis of [16] (the factor of $\gamma$ from the transformation of $\sigma_{t}$ to $\sigma_{t,0}$ can be included in the transformation of the spatial coordinate to the stretched optical depth parameter).
 Equation (63) can also be obtained from the the static boundary condition specified in comoving space-time, assuming the comoving time derivative is 0.
-Following the discretization procedure of [densmore2007], the leakage opacity and boundary transmission probability are
+Following the discretization procedure of [8], the leakage opacity and boundary transmission probability are
 $\tilde{\sigma}_{D,i\rightarrow i-1}=\frac{2}{3\gamma_{i}^{2}\Delta x_{i}(\gamma_{i-1/2}^{+}\sigma_{t,0,i-1/2}^{+}\Delta x_{i}+2\lambda)}\;\;,$ (64a)
@@
 Furthermore, $\alpha=0.5$ should furnish 2nd-order accuracy in time.
-In Section 4, we restrict to steady state for comparing accuracy of the diffusion equation to lab-frame Monte Carlo transport; thus we eliminate the issue of the time-stability considered in the present section.
+In Section 4, we consider time-independent code implementations for comparing accuracy of the diffusion equation to lab-frame Monte Carlo transport; thus we avoid the issue of the time-stability considered in the present section.
 4 Numerical results
@@
 where $d_{p}\in i$ implies the sum is over the set of tracks of particle $p$ inside cell $i$.
-We note that expanding the quadratic and distributing the sum furnishes path length estimators for lab-frame energy density, flux and pressure in the context of the standard energy-momentum tensor transformation for comoving energy density (see, for instance, [castor2004], Chapter 6).
+We note that expanding the quadratic and distributing the sum furnishes path length estimators for lab-frame energy density, flux and pressure in the context of the standard energy-momentum tensor transformation for comoving energy density (see, for instance, [5], Chapter 6).
 For the following calculations, we fix the spatial domain $x\in[0,1]$ with 128 spatial cells.
@@
 For the MC transport, we employ 512 particles per cell for the uniform volume source and 65,536 particles for the point source.
+All comparisons in this section use lab-frame time-independent code implementations of both the discrete diffusion and Monte Carlo.
 4.1 Constant $\beta$
 Here we compare the drift-diffusion equation with lab frame MC transport at several constant velocities.
+This problem has a discretized point source $Q_{0,i}=1/\Delta x$ at $x=0.5$.
 This removes the effect of the adiabatic and Doppler shift term in the drift diffusion equation, isolating the effect of the $\gamma$-factors.
@@
 The fully relativistic drift-diffusion result agrees closely at each $\beta$, but is missing a spike near the source present in the MC result.
-Figure 2: Scaled scalar flux versus spatial coordinate for several values of constant $\beta$.
+Figure 2: Scaled scalar flux versus spatial coordinate for a point source at $x=0.5$ and several values of constant $\beta$.
 The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@
 It is possible the degree of anisotropy near the higher bound in the MC solution is not captured in the asymptotic boundary condition we have implemented.
-Figure 3: Scaled scalar flux versus spatial coordinate for several values of $\Delta\beta=0.06,0.6$ in $\beta=0.6+(x-1/2)\Delta\beta$.
+Figure 3: Scaled scalar flux versus spatial coordinate for a point source at $x=0.5$ and several values of $\Delta\beta=0.06,0.6$ in $\beta=0.6+(x-1/2)\Delta\beta$.
 The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@
 Finally, we consider a problem with a very sharp, localized velocity gradient, where our velocity discretization approximates $\beta=0.3+\Theta(x-1/2)\Delta\beta$, where $\Theta(\cdot)$ is again the unit step function.
-Departing from the previous problems we consider a uniform comoving volume source.
+Departing from the previous problems we consider a uniform comoving volume source, $Q_{0}=1$.
 The jump in velocity at $x=1/2$ incurs a significant change in the comoving scalar intensity, as seen in Fig.
@@
 For the larger jump $\Delta\beta=0.6$, we see that the fully relativistic solution is in better agreement in the higher-velocity region, but all three solutions agree fairly well in the lower-velocity region.
-Figure 4: Scaled scalar flux versus spatial coordinate for several values of $\Delta\beta=0.06,0.6$ in $\beta=0.3+\Theta(x-1/2)\Delta\beta$.
+Figure 4: Scaled scalar flux versus spatial coordinate for a uniform comoving source and several values of $\Delta\beta=0.06,0.6$ in $\beta=0.3+\Theta(x-1/2)\Delta\beta$.
 The maximum value of MC transport (solid) is set to 1, and semi-relativistic (dashed) and fully relativistic diffusion (dotted) are scaled accordingly.
@@
 This analysis makes use of a family of parameter scaling that is intended to be in proximity to the standard static-material background scaling used in linear transport theory.
-The resulting fully-relativistic diffusion equation has a diffusion operator that scales as $1/\gamma^{3}$, consistent with the two-direction Poisson-Kac-based derivation of [giona2017].
+The resulting fully-relativistic diffusion equation has a diffusion operator that scales as $1/\gamma^{3}$, consistent with the two-direction Poisson-Kac-based derivation of [14].
 We have scrutinized conditions for which the asymptotic analysis gives a closed form of the diffusion equation, which are not merely the standard parameter scaling relationships in the comoving frame; these give several pathologies.
@@
     and incorporating frequency dependence.
-The standard static-material scaling rules for thermal radiative transfer are somewhat different than linear transport: the absorption opacity is scaled to be asymptotically large and one considers a matter equation with an asymptotically small heat capacity [larsen1983].
+The standard static-material scaling rules for thermal radiative transfer are somewhat different than linear transport: the absorption opacity is scaled to be asymptotically large and one considers a matter equation with an asymptotically small heat capacity [23].
 The time derivative is scaled as in the linear transport parameter scaling.
-This approach is evidently distinct from the opacity-only scaling method of [thomas1930, anderson1972] considered in the general-relativistic framework.
+This approach is evidently distinct from the opacity-only scaling method of [34, 3] considered in the general-relativistic framework.
 General relativistic codes often use a time increment that does not correspond to a simple lab-frame picture, however, and the time variable is on more equal footing with the spatial variables (see Appendix Section C).
-For 3D, we must consider the finding of [giona2017] that shows an anisotropic effective diffusion coefficient.
+For 3D, we must consider the finding of [14] that shows an anisotropic effective diffusion coefficient.
 While this result is still for discrete directions, we anticipate a similar effect for continuous direction in 3D.
@@
 $\frac{\partial E}{\partial t}+c\beta\frac{\partial E}{\partial x}-\frac{\partial}{\partial x}\left(\frac{c}{3\sigma_{t,0}}\frac{\partial E}{\partial x}\right)+\frac{4}{3}cE\frac{\partial\beta}{\partial x}+\sigma_{a,0}cE=Q_{E}-\beta Q_{M}-\frac{\partial}{\partial x}\left(\frac{1}{\sigma_{t,0}}Q_{M}\right)\;\;.$ (116)
-Equation (116) is the 1D planar form of Eq. 6.51 of [castor2004], but with the following modifications:
+Equation (116) is the 1D planar form of Eq. 6.51 of [5], but with the following modifications:
   • 1.
@@
     an additional term, $\partial(Q_{M}/\sigma_{t,0})/\partial x$, appears on the right side.
-However, Eq. 6.51 of [castor2004] is for comoving $E$ and $F$. Similarly, this is the limit we obtain of (48) ((49)) in Section 2.4, but with lab-frame $E$.
+However, Eq. 6.51 of [5] is for comoving $E$ and $F$. Similarly, this is the limit we obtain of (48) ((49)) in Section 2.4, but with lab-frame $E$.
 Appendix C The Thomas scaling in non-relativistic 1D planar geometry
-Here we briefly consider the Thomas scaling [thomas1930], as presented by [anderson1972], in a non-relativistic 1D context with constant pure absorption opacity.
+Here we briefly consider the Thomas scaling [34], as presented by [3], in a non-relativistic 1D context with constant pure absorption opacity.
 To the best of our understanding, the only small parameter is the mean-free path in these approaches.
-This should constitute a simplification of Section IV of [anderson1972].
-Using similar notation to [anderson1972], the transport equation is then
+This should constitute a simplification of Section IV of [3].
+Using similar notation to [3], the transport equation is then
 $\frac{1}{\sigma}\frac{d\psi}{d\lambda}=\frac{S}{2}-\psi\;\;,$ (117)
 where we have written the partial derivatives of the streaming operator in terms of the characteristic, or affine coordinate $\lambda$.
-Following [anderson1972], we write the asymptotic expansion directly in powers of $\sigma$.
+Following [3], we write the asymptotic expansion directly in powers of $\sigma$.
 $\psi=\sum_{k=0}^{\infty}\frac{\tilde{\psi}^{(k)}}{\sigma^{k}}\;\;.$ (118)
@@
 (25), by introducing a length scale $L$, and setting $\varepsilon=1/L\sigma$ and $\psi^{(k)}=L^{k}\tilde{\psi}^{(k)}$.
-Again following [anderson1972], matching orders to O($1/\sigma$) gives
+Again following [3], matching orders to O($1/\sigma$) gives
 $\tilde{\psi}^{(0)}=\frac{S}{2}\;\;,$ (119a)
@@
 $\frac{1}{c}\frac{\partial\phi}{\partial t}-\frac{1}{\sigma}\left(\frac{1}{c^{2}}\frac{\partial^{2}}{\partial t^{2}}+\frac{1}{3}\frac{\partial^{2}}{\partial x^{2}}\right)\phi=\sigma(S-\phi)\;\;,$ (123)
-which is nearly the Telegrapher’s equation (it is so with an imaginary relaxation coefficient of the 2nd derivative in time).
-Putting into the 4-vector notation of [anderson1972, shibata2011], Eq.
-(123) becomes
+which is nearly the Telegrapher’s equation (it is so with an imaginary relaxation coefficient of the 2nd derivative in time). Putting into the 4-vector notation of [3, 31], Eq. (123) becomes
 $u^{\alpha}\nabla_{\alpha}\phi-\frac{1}{3\sigma}\nabla_{\alpha}\left(\left(3u^{\alpha}u^{\delta}+h^{\alpha\delta}\right)\nabla_{\gamma}\phi\right)=\sigma(S-\phi)\;\;,$ (124)
-where $\alpha$ and $\delta$ here are 4-indexes, $u^{\alpha}=(1,0,0,0)$ is the 4-velocity, and $h^{\alpha\delta}=u^{\alpha}u^{\delta}-\eta^{\alpha\delta}$ is the rank-4 projection operator, as in [anderson1972, shibata2011] ($\eta^{\alpha\delta}$ here is the Minkowski metric).
-In contrast to equation 5.21 of [shibata2011], Eq.
-(124) has an additional time-like component in the 2nd derivative, which can be seen as a contribution of $\tilde{\phi}^{(1)}$ (consistent with equation 5.13 of [shibata2011]).
+where $\alpha$ and $\delta$ here are 4-indexes, $u^{\alpha}=(1,0,0,0)$ is the 4-velocity, and $h^{\alpha\delta}=u^{\alpha}u^{\delta}-\eta^{\alpha\delta}$ is the rank-4 projection operator, as in [3, 31] ($\eta^{\alpha\delta}$ here is the Minkowski metric).
+In contrast to equation 5.21 of [31], Eq.
+(124) has an additional time-like component in the 2nd derivative, which can be seen as a contribution of $\tilde{\phi}^{(1)}$ (consistent with equation 5.13 of [31]).
 Excising this time-like component would furnish a non-equilibrium diffusion equation from Eqs.